### v2

# matrix,arrows,decorations.pathmorphing

Abstract.
We show that every automorphism of the congruence completion of the extended mapping class group which preserves the set of conjugacy classes of procyclic groups generated by Dehn twists is inner and that its automorphism group is naturally isomorphic to the automorphism group of the procongruence pants complex. In the genus 00 case, we prove the stronger result that all automorphisms of the profinite completion of the extended mapping class group are inner.
AMS Math Classification: 14G32, 11R32, 14D23, 20F34, 57M10.
1. Introduction
Let $S=S_{g,n}$ be a orientable surface of genus $g(S)=g$ with $n(S)=n$ punctures. We will assume that $S$ has negative Euler characteristic: $\chi(S)=2-2g-n<0$. Let then ${\Gamma}^{\pm}(S)$ be the extended mapping class group of $S$, namely the group of isotopy classes of diffeomorphisms of $S$ and ${\Gamma}(S)$ be the mapping class group of $S$, i.e. the subgroup of ${\Gamma}^{\pm}(S)$ consisting of the mapping classes which preserve the orientation of $S$. We denote, respectively, by $\mathrm{P}\Gamma(S)$ and $\mathrm{P}\Gamma^{\pm}(S)$ the pure mapping class group and the pure extended mapping class group, namely the subgroups of ${\Gamma}(S)$ and ${\Gamma}^{\pm}(S)$ consisting of those mapping classes which fix pointwise the punctures. We also let $d(S)=3g-3+n$ be the modular dimension of $S$, that is to say the dimension of the moduli stack ${\mathcal{M}}(S)$ of smooth curves whose complex models are diffeomorphic to $S$.
Ivanov (cf. [13] and [14]) proved that all automorphisms of the extended mapping class group ${\Gamma}^{\pm}(S)$ are inner for $g(S)\geq 3$ and for $g(S)\geq 2$, $n(S)\geq 1$. McCarthy (cf. [18]) showed that this is still true, for $g(S)=2$ and $n(S)=0$, if we restrict to those automorphisms which preserve the set of conjugacy classes of Dehn twists. Korkmaz (cf. [15]) extended Ivanov’s result to the case $g(S)=1$, $n(S)\geq 3$ and $g(S)=0$, $n(S)\geq 5$. The case $g(S)=1$, $n(S)=2$ was settled by Luo in [17], where he also gave a new proof for all genera.
The proof of all the above results is based on the study of the complex of curves $C(S)$. This is the abstract simplicial complex of dimension $d(S)-1$ whose simplices are the sets of isotopy classes of essential simple closed curves on $S$ which admit disjoint representatives (such sets are called multicurves). There is a natural action of the extended mapping class ${\Gamma}^{\pm}(S)$ on $C(S)$ and the above results are obtained showing that all automorphisms of $C(S)$ are induced by this action. In fact, this action corresponds to the inner action of ${\Gamma}^{\pm}(S)$ on the set of Dehn twists, which are parametrized by the vertices of $C(S)$, and, if we denote by $\operatorname{Aut}^{\mathbb{I}}({\Gamma}^{\pm}(S))$ the group of automorphisms of ${\Gamma}^{\pm}(S)$ which preserve the set of conjugacy classes of Dehn twists, there is a series of inclusion:
$\operatorname{Inn}({\Gamma}^{\pm}(S))\subseteq\operatorname{Aut}^{\mathbb{I}}({\Gamma}^{\pm}(S))\hookrightarrow\operatorname{Aut}(C(S)),$
which are showed to be all isomorphisms for $d(S)>1$ (with the only exception of the case $g(S)=1$ and $n(S)=2$). The identity $\operatorname{Inn}({\Gamma}^{\pm}(S))=\operatorname{Aut}({\Gamma}^{\pm}(S))$ then follows from the fact that all automorphisms of ${\Gamma}^{\pm}(S)$, for $d(S)>1$, with few low genus exceptions, preserve the set of conjugacy classes of Dehn twists.
In [19], Margalit determined the automorphism group of a related complex, the so called pants graph $C_{P}(S)$. This is defined as follows. The vertices of $C_{P}(S)$ are pants decompositions (i.e. maximal multicurves) of $S$ and correspond to facets (simplices of highest dimension $d(S)-1$) of $C(S)$. Two vertices $\underline{{\alpha}}=(\alpha_{0},\alpha_{1},\ldots,\alpha_{d(S)-1})$ and $\underline{{\alpha}}^{\prime}=(\alpha^{\prime}_{0},\alpha^{\prime}_{1},\ldots,\alpha^{\prime}_{d(S)-1})$ are connected by an edge if they differ by an elementary move, that is to say: the multi-curves $\underline{{\alpha}}$ and $\underline{{\alpha}}^{\prime}$ have $d(S)-1$ elements in common, so that, up to relabelling, ${\alpha}_{i}={\alpha}^{\prime}_{i}$, for $i=1,\ldots,d(S)-1$, and the surface $S^{\prime}$ obtained cutting $S$ along the curves ${\alpha}_{i}$, for $i>0$, is a surface of modular dimension $1$, i.e. $S^{\prime}=S_{1,1}$ or $S^{\prime}=S_{0,4}$. Then, ${\alpha}_{0}$ and ${\alpha}^{\prime}_{0}$, which are supported on $S^{\prime}$, should intersect in a minimal way, i.e. they have geometric intersection number $1$, in the first case, and $2$, in the second case.
Margalit’s results (cf. Theorem 1 and 2 in [19]) then imply that, for $d(S)>1$, the natural action of ${\Gamma}^{\pm}(S)$ on $C_{P}(S)$ induces a series of isomorphisms:
(1) $\operatorname{Inn}({\Gamma}^{\pm}(S))\cong\operatorname{Aut}^{\mathbb{I}}({\Gamma}^{\pm}(S))\cong\operatorname{Aut}(C_{P}(S)).$
In the paper [6], we proved a partial analogue of the above series of isomorphisms in the setting of procongruence mapping class groups which we now proceed to define.
The profinite mapping class groups ${\widehat{\Gamma}}^{\pm}(S)$, ${\widehat{\Gamma}}(S)$, $\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)$ and $\mathrm{P}{\widehat{\Gamma}}(S)$ are defined to be the profinite completions of ${\Gamma}^{\pm}(S)$, ${\Gamma}(S)$, $\mathrm{P}\Gamma^{\pm}(S)$ and $\mathrm{P}\Gamma(S)$, respectively. The procongruence mapping class groups ${\check{\Gamma}}^{\pm}(S)$, ${\check{\Gamma}}(S)$, $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ and $\mathrm{P}{\check{\Gamma}}(S)$ are the images of ${\widehat{\Gamma}}^{\pm}(S)$, ${\widehat{\Gamma}}(S)$, $\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)$ and $\mathrm{P}{\widehat{\Gamma}}(S)$, respectively, in the profinite group $\operatorname{Out}(\widehat{\pi_{1}(S)})$, where, for an abstract group $G$, we denote by $\widehat{G}$ its profinite completion. By a classical result of Grossman, the natural homomorphism from each of the above mapping class group to either its profinite or procongruence completion is injective and we then identify the abstract groups with their images in the corresponding profinite groups. By the results of [3] and [5], the profinite and procongruence completion of mapping class groups coincide if $g(S)\leq 2$. For $g(S)\geq 3$, this is an open problem. We will then rather stick with the procongruence completion, since, in contrast with the profinite completion, some basic combinatorial properties are known, thanks to the results contained in [4] and [5].
In analogy with the topological case, for the study of the automorphism group of procongruence mapping class groups, it is useful to introduce the procongruence curve complex ${\check{C}}(S)$. This is an abstract simplicial profinite complex (cf. Definition 3.2 in [4]) of dimension $d(S)-1$, naturally associated to the congruence completion of the mapping class group and endowed with a natural continuous action of ${\check{\Gamma}}(S)$. Roughly speaking (for the precise definition we refer the reader to Section 4 of [4] or Section 4.6 in [6]), ${\check{C}}(S)$ is obtained as the inverse limit of the quotients of $C(S)$ by the action of congruence levels of ${\Gamma}(S)$, that is to say finite index subgroups of ${\Gamma}(S)$ which are open for the congruence topology.
The set of profinite Dehn twists of $\mathrm{P}{\check{\Gamma}}(S)$ is defined to be the closure, inside this group, of the set of Dehn twists of $\mathrm{P}\Gamma(S)\subset\mathrm{P}{\check{\Gamma}}(S)$. The key property of ${\check{C}}(S)$ is then that its set of $k$-simplices parameterizes the profinite set $\{{\hat{\mathrm{I}}}_{\sigma}|\,{\sigma}\in{\check{C}}(S)_{k}\}$ of closed abelian subgroups of $\mathrm{P}{\check{\Gamma}}(S)$ of rank $k+1$, topologically generated by the profinite Dehn twists of $\mathrm{P}{\check{\Gamma}}(S)$ (cf. Theorem 6.9 in [4]). Here, for ${\sigma}\in{\check{C}}(S)$, we let ${\hat{\mathrm{I}}}_{\sigma}$ be the closed abelian subgroup of $\mathrm{P}{\check{\Gamma}}(S)$ topologically generated by the profinite Dehn twists parameterized by the vertices of ${\sigma}$.
The natural action of ${\check{\Gamma}}^{\pm}(S)$ on ${\check{C}}(S)$ then corresponds to the conjugacy action of ${\check{\Gamma}}^{\pm}(S)$ on the set of abelian groups parameterized by ${\check{C}}(S)$ and, if we denote by $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$ (resp. $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$) the group of those automorphisms of ${\check{\Gamma}}^{\pm}(S)$ (resp. $\mathrm{P}{\check{\Gamma}}(S)$) which preserve the set of conjugacy classes of the procyclic subgroups generated by profinite Dehn twists, we get a series of natural monomorphism:
$\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\subseteq\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\hookrightarrow\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))\hookrightarrow\operatorname{Aut}({\check{C}}(S)).$
Unlike in the topological case, however, this is not going to be a series of isomorphisms. A procongruence analogue of the pants graph $C_{P}(S)$ turns out to be more useful here.
The procongruence pants complex ${\check{C}}_{P}(S)$ is the profinite graph which is, roughly speaking (cf. Section 6.2 in [6] for the precise definition), the inverse limit of the quotients of $C_{P}(S)$ by the action of congruence levels of ${\Gamma}(S)$. It is also endowed, by definition, with a natural continuous action of ${\check{\Gamma}}^{\pm}(S)$. The profinite set of vertices of ${\check{C}}_{P}(S)$ identifies with the profinite set of facets of ${\check{C}}(S)$ and so it parameterizes the profinite set $\{{\hat{\mathrm{I}}}_{\sigma}|\,{\sigma}\in{\check{C}}(S)_{d(S)-1}\}$ of maximal abelian subgroups of $\mathrm{P}{\check{\Gamma}}(S)$ topologically generated by profinite Dehn twists. The natural continuous action of ${\check{\Gamma}}^{\pm}(S)$ on ${\check{C}}_{P}(S)$ is then induced by the conjugacy action of ${\check{\Gamma}}^{\pm}(S)$ on this profinite set.
The main result of the paper is an analogue of the series of isomorphisms (1):
Theorem 1.1.
For a connected hyperbolic surface $S$ such that $d(S)>1$, there is a series of natural isomorphisms:
$\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\cong\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))\cong\operatorname{Aut}({\check{C}}_{P}(S)).$
Remark 1.2.
The condition on automorphisms in Theorem 1.1 is slightly more restrictive than the one considered in [6], where we denoted by $\operatorname{Aut}^{\ast}({\check{\Gamma}}(S))$ the subgroup of $\operatorname{Aut}({\check{\Gamma}}(S))$ consisting of those automorphisms which preserve the conjugacy classes of decomposition subgroups of ${\check{\Gamma}}(S)$ (cf. Definition 7.1 in [6]). It is not difficult to see that there is an inclusion $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}(S))\subseteq\operatorname{Aut}^{\ast}({\check{\Gamma}}(S))$ which for $d(S)>1$ and $Z({\check{\Gamma}}(S))=\{1\}$ is indeed an equality but otherwise is strict.
In [6], we defined the arithmetic procongruence mapping class group ${\check{\Gamma}}_{\mathbb{Q}}(S)$ to be the image of the étale fundamental group of the moduli stack ${\mathcal{M}}(S)_{\mathbb{Q}}:={\mathcal{M}}(S)\times\operatorname{Spec}(\mathbb{Q})$ by the monodromy representation associated to the universal curve over ${\mathcal{M}}(S)_{\mathbb{Q}}$. We then proved that, for $d(S)>1$, there holds (cf. Theorem 9.16 in [6]):
$\operatorname{Inn}({\check{\Gamma}}_{\mathbb{Q}}(S))=\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}_{\mathbb{Q}}(S)).$
The identity $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$ in Theorem 1.1 is the real analogue of this result. The extended procongruence mapping class group ${\check{\Gamma}}^{\pm}(S)$ can indeed be defined to be the image of the étale fundamental group of the moduli stack ${\mathcal{M}}(S)_{\mathbb{R}}:={\mathcal{M}}(S)\times\operatorname{Spec}(\mathbb{R})$ by the monodromy representation associated to the universal curve over ${\mathcal{M}}(S)_{\mathbb{R}}$.
Theorem 1.1 is more striking than Theorem 9.16 in [6], since ${\check{\Gamma}}^{\pm}(S)$ contains ${\check{\Gamma}}(S)$ as a normal index $2$ subgroup, rather than as a normal subgroup of infinite index, and we know that $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}(S))$ is quite far from consisting only of inner automorphisms. In fact, the outer automorphism group $\operatorname{Out}^{\mathbb{I}}({\check{\Gamma}}(S))$ contains a copy of the absolute Galois group of the rationals $G_{\mathbb{Q}}$ (cf. Corollary 7.6 in [4]).
The analogy between Theorem 1.1 and Margalit’s series of isomorphisms (1) falls short only in that, in contrast with the topological case, we do not know whether, for $d(S)>1$ and $Z({\check{\Gamma}}(S))=\{1\}$, there holds $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))=\operatorname{Aut}({\check{\Gamma}}^{\pm}(S))$. However, thanks to a recent result by Hoshi, Minamide and Mochizuki (cf. [11]), we are able to fill this gap in genus 00. Since, in this case, we also know that the congruence subgroup property holds, we get:
Theorem 1.3.
For $g(S)=0$ and $n(S)\geq 5$, there holds:
$\operatorname{Inn}({\widehat{\Gamma}}^{\pm}(S))=\operatorname{Aut}({\widehat{\Gamma}}^{\pm}(S))=\operatorname{Aut}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)).$
This result provides, to our knowledge, the first example of a finitely generated, infinite, residually finite, complete group whose profinite completion is also complete.
Theorem 1.3 can also be rephrased as a real anabelian property for the moduli spaces ${\mathcal{M}}_{0,n}$ of $n$-pointed, genus zero curves. Let $({\mathcal{M}}_{0,n})_{\mathbb{R}}:={\mathcal{M}}_{0,n}\times\operatorname{Spec}\mathbb{R}$, let $\pi_{1}^{\mathrm{et}}(({\mathcal{M}}_{0,n})_{\mathbb{R}})$ be its étale fundamental group for some choice of geometric base point and let $\Sigma_{n}$ be the symmetric group on $n$ letters. We then have:
Corollary 1.4.
For $n\geq 5$, there is a natural isomorphism:
$\operatorname{Aut}(({\mathcal{M}}_{0,n})_{\mathbb{R}})\cong\operatorname{Out}(\pi_{1}^{\mathrm{et}}(({\mathcal{M}}_{0,n})_{\mathbb{R}}))\cong\Sigma_{n}.$
Another interesting application of Theorem 1.1 is the following. Since $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))$ identifies with a subgroup of $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}(S))$ and $G_{\mathbb{Q}}$ embeds in $\operatorname{Out}^{\mathbb{I}}({\check{\Gamma}}(S))$, we have (cf. (ii) of Proposition 4 in [16], for the genus 00 case):
Corollary 1.5.
For a connected hyperbolic surface $S$ such that $d(S)>1$, the subgroup $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))$ is its own normalizer in $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}(S))$. In particular, the image of an element of $G_{\mathbb{Q}}$ corresponding to complex conjugation is self-centralizing in $\operatorname{Out}^{\mathbb{I}}({\check{\Gamma}}(S))$.
A few words about the proof of Theorem 1.1. The isomorphism $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\cong\operatorname{Aut}({\check{C}}_{P}(S))$ is a refinement of Theorem 8.1 in [6]. Here, we are able to show that there is indeed a coherent and symmetric way to define an orientation on the procongruence pants complex ${\check{C}}_{P}(S)$ (cf. Section 2.2). The proof of the identity $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ is then based on a further improvement of this isomorphism. An element $f$ of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ acts on the vertex set of ${\check{C}}_{P}(S)$, since this is naturally identified with the profinite set $\{{\hat{\mathrm{I}}}_{\sigma}|\,{\sigma}\in{\check{C}}(S)_{d(S)-1}\}$. Theorem 2.15 then states that $f$ is induced by an inner automorphism of ${\check{\Gamma}}^{\pm}(S)$ as soon as this action sends the vertex set of an edge of ${\check{C}}_{P}(S)$ to the vertex set of another edge. In order to prove the identity $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$, we then have to show that the elements of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))\subset\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ have this property. By an induction argument, it is actually enough to show that this is true for $g(S)=0$.
This is accomplished by considering the action of antiholomorphic involutions of $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ on the procongruence curve complex ${\check{C}}(S)$. By Lemma 4.2, for $g(S)=0$, there is only one ${\check{\Gamma}}^{\pm}(S)$-conjugacy class of antiholomorphic involutions in $\mathrm{P}{\check{\Gamma}}(S)$. We can then assume that a given automorphism of $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ fixes this involution and then preserves its fixed point locus in ${\check{C}}(S)$. By Lemma 4.3 and Proposition 3.6, this locus is finite and consists of isotopy classes of simple closed curves on $S$ which have between them geometric intersection either 00 or $2$. Since pairs of curves with geometric intersection $2$ correspond to edges of the pants complex, we can apply Theorem 2.15 and conclude that the given automorphism is induced by an inner automorphism of ${\check{\Gamma}}^{\pm}(S)$.
Acknowledgements. We thank Javier Aramayona and Dan Margalit for some useful comments on a preliminary version of this manuscript which helped us to spot an error in the proof and improve the main result of the paper.
2. Two preliminary results
2.1. Augmented Teichmüller spaces and procongruence moduli stacks of curves
Let ${\mathcal{T}}(S)$ be the Teichmüller space associated to the surface $S$ and $\overline{{\mathcal{T}}}(S)$ be the augmented Teichmüller space (in [6], we called it the Bers bordification of ${\mathcal{T}}(S)$). The latter can be defined as the completion of the Teichmüller space ${\mathcal{T}}(S)$ with respect to the Weil-Petersson metric (cf. Theorem 4 in [26]). The augmented Teichmüller space $\overline{{\mathcal{T}}}(S)$ is a partial ${\Gamma}^{\pm}(S)$-equivariant compactification of ${\mathcal{T}}(S)$ such that the quotient $\overline{{\mathcal{T}}}(S)/{\Gamma}(S)$ is isomorphic to the coarse moduli space $\overline{M}(S)$ of the the DM compactification ¯${\overline{\mathcal{M}}}(S)$ of ${\mathcal{M}}(S)$.
For $S$ a disconnected hyperbolic surface, let ${\mathcal{T}}(S)$ (resp. $\overline{{\mathcal{T}}}(S)$) be the direct product of the Teichmüller spaces (resp. augmented Teichmüller spaces) associated to the connected components of the surface $S$. The closed strata of codimension $k+1$ in $\overline{{\mathcal{T}}}(S)$ of the boundary $\partial\overline{{\mathcal{T}}}(S):=\overline{{\mathcal{T}}}(S)\smallsetminus{\mathcal{T}}(S)$ are then parameterized by the $k$-simplices of $C(S)$, and, for ${\sigma}\in C(S)_{k}$, there is a natural isomorphism $\partial\overline{{\mathcal{T}}}(S)_{\sigma}\cong\overline{{\mathcal{T}}}(S\smallsetminus{\sigma})$, where we denote by $\partial\overline{{\mathcal{T}}}(S)_{\sigma}$ the closed stratum associated to ${\sigma}$.
Let us denote by ~$\widetilde{{\mathcal{F}}}(S)$ the $1$-dimensional stratum of the boundary $\partial\overline{{\mathcal{T}}}(S)$. Then, each irreducible component of ~$\widetilde{{\mathcal{F}}}(S)$ is isomorphic to the cuspidal bordification ¯$\overline{{\mathbb{H}}}={\mathbb{H}}\cup{\mathbb{P}}^{1}(\mathbb{Q})$ of the hyperbolic plane ${\mathbb{H}}$ and $C_{P}(S)$ identifies with the $1$-skeleton of a ${\Gamma}(S)$-equivariant triangulation of ~$\widetilde{{\mathcal{F}}}(S)$.
[…]
2.2. Automorphisms of the procongruence pants complex and orientations
For $d(S)=1$, the pants complex $C_{P}(S)$ coincides with the Farey graph $\rm F$. This is the $1$-skeleton of a $2$-dimensional simplicial complex $\Delta$, whose geometric realization $|\Delta|$ identifies with a tessellation of the cuspidal bordification $\overline{{\mathbb{H}}}$ of the hyperbolic plane. The homeomorphism $|\Delta|\cong\overline{{\mathbb{H}}}$ becomes a conformal isomorphism when $|\Delta|$ is given the piecewise equilateral flat structure. The orientation of the Farey graph $\rm F$ (and so of $C_{P}(S)$ for $d(S)=1$) is then simply the orientation of $\Delta$ associated to the complex structure of $\overline{{\mathbb{H}}}$.
[…]
For a multicurve ${\sigma}$, let $C_{P}(S\smallsetminus{\sigma})$ be the disjoint union of the pants complexes associated to the connected components of $S\smallsetminus{\sigma}$ and let $d(S\smallsetminus{\sigma})$ be the sum of the modular dimensions of the connected components of $S\smallsetminus{\sigma}$. For ${\sigma}\in C(S)_{d(S)-2}$, we have that $d(S\smallsetminus{\sigma})=1$, so that the pants complex $C_{P}(S\smallsetminus{\sigma})$ is isomorphic to the Farey graph $\rm F$ and identifies with a subgraph of $C_{P}(S)$, which we denote by $\rm F_{\sigma}$. The pants complex $C_{P}(S)$ is then the infinite union of the Farey subgraphs $\{\rm F_{\sigma}\}_{{\sigma}\in C(S)_{d(S)-2}}$ and we give each of them the orientation defined above. A corollary of Margalit’s series of isomorphisms (1) is then that automorphisms of $C_{P}(S)$ either preserve or reverse all orientations of the Farey subgraphs.
The procongruence pants complex ${\check{C}}_{P}(S)$, for $d(S)>1$ is the union of the profinite set of profinite Farey graphs $\{{\widehat{\operatorname{F}}}_{\sigma}\}_{{\sigma}\in{\check{C}}(S)_{d(S)-2}}$, each naturally associated to a $(d(S)-2)$-simplex of the procongruence curve complex ${\check{C}}(S)$ (cf. Definition 6.3 in [6]), and we give each of them the orientation defined above. However, it is not clear that the automorphisms of ${\check{C}}_{P}(S)$ act in synchrony on the orientations of its profinite Farey subgraphs.
Let us denote by $O(S)$ the finite set of the topological types of $(d(S)-1)$-multicurves on $S$. To remedy the above issue, in [6], we associated a character $\operatorname{Aut}({\check{C}}_{P}(S))\to\{\pm 1\}$ to each ${\sigma}\in O(S)$ in the following way. The tautological action of $\operatorname{Aut}({\check{C}}_{P}(S))$ on ${\check{C}}_{P}(S)$ preserves profinite Farey subgraphs and ${\check{\Gamma}}(S)$-orbits of profinite Farey subgraphs. We assigned to an automorphism $\phi\in\operatorname{Aut}({\check{C}}_{P}(S))$ the plus or minus sign according to whether $\phi$ sends or not the fixed orientation of ${\widehat{\operatorname{F}}}_{\sigma}$ to the orientation of $\phi({\widehat{\operatorname{F}}}_{\sigma})$. We then proved that there is an exact sequence, for $d(S)>1$ and $S\neq S_{1,2}$ (cf. Theorem 8.1 in [6]):
(2) $1\to\operatorname{Inn}({\check{\Gamma}}(S))\to\operatorname{Aut}({\check{C}}_{P}(S))\to\prod_{O(S)}\{\pm 1\}.$
For $S=S_{1,2}$, the group $\operatorname{Aut}({\check{C}}_{P}(S_{1,2}))$ must be replaced with the subgroup of those automorphisms preserving the set of separating curves.
The first result of this section is an improved version of Theorem 8.1 in [6]. We will show that all the above characters are in fact synchronized and that the case $S\neq S_{1,2}$ is not exceptional, so that we have:
Proposition 2.1.
For a connected hyperbolic surface $S$ such that $d(S)>1$, there is a short exact sequence:
$1\to\operatorname{Inn}({\check{\Gamma}}(S))\to\operatorname{Aut}({\check{C}}_{P}(S))\to\{\pm 1\}\to 1$
2.3. The case $S=S_{0,5}$
There is only one topological type of 00-simplices in $C(S_{0,5})$. Hence, the exact sequence (2) takes the simple form:
$1\to\operatorname{Inn}({\check{\Gamma}}(S_{0,5}))\to\operatorname{Aut}({\check{C}}_{P}(S_{0,5}))\to\{\pm 1\}.$
That this sequence is also right exact follows considering the action on ${\check{C}}_{P}(S_{0,5})$ of an inner automorphism $\operatorname{inn}f$ for $f\in{\Gamma}^{\pm}(S_{0,5})\smallsetminus{\Gamma}(S_{0,5})$. This proves Proposition 2.1 for $S=S_{0,5}$. Let us make some additional remarks which will be useful for the case $S=S_{1,2}$.
Let $\partial\overline{{\mathbb{M}}}(S_{0,5})$ be the DM boundary of the inverse limit $\overline{{\mathbb{M}}}(S_{0,5})$. As explained in Secion 2.1, the profinite pants complex ${\check{C}}_{P}(S_{0,5})$ identifies with the $1$-skeleton of a triangulation of $\partial\overline{{\mathbb{M}}}(S_{0,5})$. Let $\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$ be the group of automorphism which restrict to a conformal or an anticonformal map on each irreducible component and $\operatorname{Aut}^{+}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$ its subgroup consisting of conformal automorphisms.
Lemma 2.2.
The natural action of ${\check{\Gamma}}(S_{0,5})$ on $\partial\overline{{\mathbb{M}}}(S_{0,5})$ induces an isomorphism ${\check{\Gamma}}(S_{0,5})\cong\operatorname{Aut}^{+}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$.
Proof.
This follows from Lemma 8.8 in [6]. ∎
Lemma 2.3.
$\operatorname{Aut}^{+}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$ is an index $2$ subgroup of $\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$, that is to say, every automorphism of $\partial\overline{{\mathbb{M}}}(S_{0,5})$ either preserves or reverses the orientation of all irreducible components simultaneously. In particular, there is a natural isomorphism ${\check{\Gamma}}^{\pm}(S_{0,5})\cong\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$.
Proof.
The irreducible components of $\partial\overline{{\mathbb{M}}}(S_{0,5})$ are parameterized by the 00-simplices in ${\check{C}}(S_{0,5})$. Let us denote by $\partial\overline{{\mathbb{M}}}(S_{0,5})_{\sigma}$ the irreducible component associated to the 00-simplex ${\sigma}\in{\check{C}}(S_{0,5})_{0}$. Let us then define the character $\chi_{\sigma}\colon\thinspace\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))\to\{\pm 1\}$ which takes the value $+1$ on $f\in\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$ if $f$ sends the standard orientation of $\partial\overline{{\mathbb{M}}}(S_{0,5})_{\sigma}$ to the standard orientation of $f(\partial\overline{{\mathbb{M}}}(S_{0,5})_{\sigma})$ and $-1$ otherwise. There is an exact sequence:
$1\to\operatorname{Aut}^{+}(\partial\overline{{\mathbb{M}}}(S_{0,5}))\to\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))\to\prod_{{\sigma}\in{\check{C}}(S_{0,5})_{0}}\{\pm 1\}.$
In particular, $\operatorname{Aut}^{+}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$ is a normal subgroup of $\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$.
We can now argue exactly as in Section 8.7 of [6] and conclude that the representation $\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))\to\prod_{{\sigma}\in{\check{C}}(S_{0,5})_{0}}\{\pm 1\}$ is constant on the ${\check{\Gamma}}(S_{0,5})$-orbit of 00-simplices of ${\check{C}}(S_{0,5})$. Since there is only one such orbit, the first statement of the lemma follows. The second statement then follows from Lemma 2.2. ∎
Lemma 2.4.
There is a natural isomorphism $\operatorname{Aut}({\check{C}}_{P}(S_{0,5}))\cong\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$.
Proof.
By Proposition 8.3 in [6], there is a natural monomorphism $\operatorname{Aut}({\check{C}}_{P}(S_{0,5}))\hookrightarrow\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$. The conclusion then follows from Lemma 2.3 and the fact that ${\check{\Gamma}}^{\pm}(S_{0,5})$ identifies with a subgroup of $\operatorname{Aut}({\check{C}}_{P}(S_{0,5}))$. ∎
2.4. Proof of Proposition 2.1 for $S=S_{1,2}$
For $[C,P_{1},P_{2}]\in{\mathcal{M}}_{1,[2]}$, there is a unique elliptic involution ${\upsilon}$ on $C$ and, if we denote by $C_{\upsilon}$ the (genus 00) quotient of $C$ by this involution and by $B_{\upsilon}$ the branch locus of the orbit map $C\to C_{\upsilon}$, then, the assignment $[C,P_{1},P_{2}]\mapsto[C_{\upsilon},Q,B_{\upsilon}]$, where $Q$ is the image of the pair of points $P_{1},P_{2}$ in $C_{\upsilon}$, defines a morphism of DM stacks ${\mathcal{M}}_{1,[2]}\to{\mathcal{M}}_{0,1[4]}$, where we denote by ${\mathcal{M}}_{0,1[4]}$ the moduli stack of genus 00 projective smooth curves with $5$ labeled points, one of which is singled out and the others are left unordered. The morphism ${\mathcal{M}}_{1,[2]}\to{\mathcal{M}}_{0,1[4]}$ is a $\mathbb{Z}/2$-gerbe which is split, since the composition ${\mathcal{M}}_{1,2}\to{\mathcal{M}}_{1,[2]}\to{\mathcal{M}}_{0,1[4]}$ is an isomorphism of DM stacks.
The $\mathbb{Z}/2$-gerbe ${\mathcal{M}}_{1,[2]}\to{\mathcal{M}}_{0,1[4]}$ then induces on topological fundamental groups a split short exact sequence:
$1\to\langle{\upsilon}\rangle\to\pi_{1}({\mathcal{M}}_{1,[2]})\to\pi_{1}({\mathcal{M}}_{0,1[4]})\to 1$
In terms of mapping class groups, this can be described as follows. Let ${\upsilon}\in{\Gamma}(S_{1,2})$ be the hyperelliptic involution, let $S_{/{\upsilon}}$ be the quotient of the surface $S_{1,2}$ by ${\upsilon}$ and let $B_{\upsilon}$ be the branch locus of the orbit map $S\to S_{/{\upsilon}}$. The surface $S_{/{\upsilon}}$ is a $1$-punctured sphere, there is a diffeomorphism $S_{/{\upsilon}}\smallsetminus B_{\upsilon}\cong S_{0,5}$ and, if we denote by $Q$ the puncture of $S_{0,5}$ which corresponds to the puncture of $S_{/{\upsilon}}$ via the above diffeomorphism, by Birman-Hilden theory, there is a short exact sequence (cf. Theorem 2.3 in [8]):
$1\to\langle{\upsilon}\rangle\to{\Gamma}(S_{1,2})\to{\Gamma}(S_{0,5})_{Q}\to 1,$
where ${\Gamma}(S_{0,5})_{Q}$ is the stabilizer of the puncture $Q$ in ${\Gamma}(S_{0,5})$. There is then a natural isomorphism between this short exact sequence and the one obtained above in a geometric way. In particular, there is a natural isomorphism $\mathrm{P}\Gamma(S_{1,2})\cong{\Gamma}(S_{0,5})_{Q}$ so that $\mathrm{P}\Gamma(S_{0,5})$ identifies with a subgroup of $\mathrm{P}\Gamma(S_{1,2})$. We record the following for future use:
Lemma 2.5.
The group $\mathrm{P}\Gamma(S_{0,5})$ identifies with the normal subgroup of $\mathrm{P}\Gamma(S_{1,2})$ generated by squares of nonseparating Dehn twists.
Proof.
The image of a nonseparating Dehn twist via the epimorphism ${\Gamma}(S_{1,2})\to{\Gamma}(S_{0,5})_{Q}$ is a braid twist. Hence, the image of the square of a nonseparating Dehn twist is a Dehn twist about a simple closed curve on $S_{0,5}$ which bounds a $2$-punctured disc. The subgroup of ${\Gamma}(S_{1,2})$ generated by squares of nonseparating Dehn twists is contained in $\mathrm{P}\Gamma(S_{1,2})$ and has trivial intersection with $\langle{\upsilon}\rangle$. Hence, it identifies with the subgroup of ${\Gamma}(S_{0,5})_{Q}$ generated by Dehn twists about simple closed curves on $S_{0,5}$ bounding $2$-punctured discs, which is $\mathrm{P}\Gamma(S_{0,5})$. ∎
From the above discussion, we see, in particular, that there is a natural finite étale morphism ${\mathcal{M}}_{0,5}\to{\mathcal{M}}_{1,2}$, so that ${\mathcal{M}}_{0,5}$ identifies with a level structure over ${\mathcal{M}}_{1,[2]}$. Moreover, the DM compactification ¯${\overline{\mathcal{M}}}_{0,5}$ of ${\mathcal{M}}_{0,5}$ coincides with the DM compactification of the latter as a level structure. Therefore, there are natural isomorphisms ${\mathbb{M}}(S_{0,5})\cong{\mathbb{M}}(S_{1,2})$, $\overline{{\mathbb{M}}}(S_{0,5})\cong\overline{{\mathbb{M}}}(S_{1,2})$ and then $\partial\overline{{\mathbb{M}}}(S_{0,5})\cong\partial\overline{{\mathbb{M}}}(S_{1,2})$.
The Teichmüller spaces ${\mathcal{T}}(S_{0,5})$ and ${\mathcal{T}}(S_{1,2})$ are the universal covers of ${\mathcal{M}}_{0,5}$ and ${\mathcal{M}}_{1,2}$, respectively. Hence, there is a natural ${\Gamma}(S_{1,2})$-equivariant isomorphism ${\mathcal{T}}(S_{0,5})\cong{\mathcal{T}}(S_{1,2})$. There is also a compatible ${\Gamma}(S_{1,2})$-equivariant isomorphism of curve complexes $C(S_{0,5})\cong C(S_{1,2})$. Since the Weil-Petersson metric on the Teichmüller space is determined, modulo strong equivalence, by the Fenchel-Nielsen coordinates, it follows that the isomorphism ${\mathcal{T}}(S_{0,5})\cong{\mathcal{T}}(S_{1,2})$ induces a natural ${\Gamma}(S_{1,2})$-equivariant isomorphism of augmented Teichmüller spaces $\overline{{\mathcal{T}}}(S_{0,5})\cong\overline{{\mathcal{T}}}(S_{1,2})$ and then $\partial\overline{{\mathcal{T}}}(S_{0,5})\cong\partial\overline{{\mathcal{T}}}(S_{1,2})$.
Even though the pants graph $C_{P}(S_{1,2})$ is the $1$-skeleton of a ${\Gamma}(S_{1,2})$-equivariant triangulation of $\partial\overline{{\mathcal{T}}}(S_{1,2})$, the same holds for $C_{P}(S_{0,5})$ and $\partial\overline{{\mathcal{T}}}(S_{0,5})$ and the vertex sets of the pants graphs $C_{P}(S_{1,2})$ and $C_{P}(S_{0,5})$ naturally identify, there is no natural map between these two graphs to account for it. This follows from a general result by Aramayona (cf. Theorem A in [1]) but also from a careful analysis of the pairs of curves with minimal intersection occurring in $S_{1,2}$ and $S_{0,5}$.
In any case, from the natural ${\Gamma}(S_{1,2})$-equivariant bijective map of 00-simplices $C_{P}(S_{1,2})_{0}\cong C_{P}(S_{0,5})_{0}$, passing to ${\check{\Gamma}}(S_{1,2})$-completions, we get a natural ${\check{\Gamma}}(S_{1,2})$-equivariant bijective map ${\check{C}}_{P}(S_{1,2})_{0}\cong{\check{C}}_{P}(S_{0,5})_{0}$. By Proposition 8.3 in [6], there is a natural monomorphism $\operatorname{Aut}({\check{C}}_{P}(S_{1,2}))\hookrightarrow\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{1,2}))$ and then $\operatorname{Aut}({\check{C}}_{P}(S_{1,2}))\hookrightarrow\operatorname{Aut}(\partial\overline{{\mathbb{M}}}(S_{0,5}))$. From Lemma 2.4, it then follows that every continuous automorphism of ${\check{C}}_{P}(S_{1,2})$ induces one of ${\check{C}}_{P}(S_{0,5})$ (compatible on vertex sets with the ${\check{\Gamma}}(S_{1,2})$-equivariant bijection given above).
For a given $f\in\operatorname{Aut}({\check{C}}_{P}(S_{1,2}))$, let us denote by $\tilde{f}$ the induced automorphism of ${\check{C}}_{P}(S_{0,5})$. By Lemma 2.3, $\operatorname{Aut}(C_{P}(S_{0,5}))$ identifies with a dense subgroup of the profinite group $\operatorname{Aut}({\check{C}}_{P}(S_{0,5}))$ and $\operatorname{Inn}({\check{\Gamma}}(S_{1,2}))$ with an open subgroup of the same group. Therefore, after possibly composing $\tilde{f}$ with an element of $\operatorname{Inn}({\check{\Gamma}}(S_{1,2}))$, we can assume that $\tilde{f}\in\operatorname{Aut}(C_{P}(S_{0,5}))$. In particular, the given $f$ preserves the vertex set $C_{P}(S_{1,2})_{0}\subset{\check{C}}_{P}(S_{1,2})_{0}$.
Lemma 2.6.
If the vertices of an edge $e$ of the procongruence pants graph ${\check{C}}_{P}(S)$ belong to $C_{P}(S)_{0}\subset{\check{C}}_{P}(S)_{0}$, then $e\in C_{P}(S)_{1}\subset{\check{C}}_{P}(S)_{1}$.
Proof.
Let $\{{\sigma}_{0},{\sigma}_{1}\}\subset C(S)_{d(S)-1}\subset{\check{C}}(S)_{d(S)-1}$ be the vertex set of $e$. By the first item of Lemma 6.6 in [6], the edge $e$ is then contained in the profinite Farey subgraph ${\widehat{\operatorname{F}}}_{{\sigma}_{0}\cap{\sigma}_{1}}$, which is obtained as the $\widehat{\operatorname{PSL}_{2}(\mathbb{Z})}$-completion of the Farey subgraph $\rm F_{{\sigma}_{0}\cap{\sigma}_{1}}\subset C_{P}(S)$. Hence it is enough to prove the statement of the lemma for the profinite Farey graph ${\widehat{\operatorname{F}}}$.
Given two vertices $v_{0},v_{1}\in\rm F_{0}$, there is a unique geodesic ${\gamma}$ in $\overline{{\mathbb{H}}}$ connecting these two points and a finite index subgroup ${\Gamma}^{\lambda}$ of $\operatorname{PSL}_{2}(\mathbb{Z})$ such that, for all ${\Gamma}^{{\lambda}^{\prime}}\leq{\Gamma}^{\lambda}$, the image of ${\gamma}$ in the quotient surface ¯$\overline{{\mathbb{H}}}/{\Gamma}^{{\lambda}^{\prime}}$ is a simple geodesic arc. This implies that, if $\{v_{0},v_{1}\}$ is the vertex set of an edge of ${\widehat{\operatorname{F}}}$, the distance between $v_{0}$ and $v_{1}$ in $\overline{{\mathbb{H}}}$ is $1$, which is possible only if $\{v_{0},v_{1}\}$ is the vertex set of an edge of $\rm F$. ∎
From Lemma 2.6, it follows that $f$ induces an automorphism of the pants complex $C_{P}(S_{1,2})$. By Theorem 1 and Theorem 2 of [19], we then have that $f\in\operatorname{Inn}({\Gamma}^{\pm}(S_{1,2}))$ which completes the proof of the case $S=S_{1,2}$ of Proposition 2.1.
2.5. Four lemmas
The following definition will play a fundamental role in the proof, by induction, of the general case of Proposition 2.1:
Definition 2.7.
For $d(S)>1$, every $(d(S)-2)$-multicurve on $S$ contains at least a simple closed curve which is either nonseparating or bounds a $2$-punctured disc. For a fixed such simple closed curve ${\gamma}$, we then let $L_{\gamma}$ be the closed subgraph of ${\check{C}}_{P}(S)$ which is the union of all profinite Farey subgraphs ${\widehat{\operatorname{F}}}_{\sigma}$ such that ${\gamma}\in{\sigma}$.
Let us denote by $S_{\gamma}$ either $S\smallsetminus{\gamma}$, for ${\gamma}$ nonseparating, or the connected component of $S\smallsetminus{\gamma}$ of positive modular dimension, for ${\gamma}$ bounding a $2$-punctured disc. We then have:
Lemma 2.8.
The profinite subgraph $L_{\gamma}$ of ${\check{C}}_{P}(S)$ is naturally isomorphic to the procongruence pants complex ${\check{C}}_{P}(S_{\gamma})$.
Proof.
By Remark 4.7 in [4], the link $\operatorname{Lk}({\gamma})\subset{\check{C}}(S)$ is naturally isomorphic to ${\check{C}}(S_{\gamma})$. Let $\xi\colon\thinspace{\check{C}}(S_{\gamma})\stackrel{{\scriptstyle\sim}}{{\to}}\operatorname{Lk}({\gamma})$ be such isomorphism. We can then identify the vertex set of ${\check{C}}_{P}(S_{\gamma})$ with a subset of the vertex set of ${\check{C}}_{P}(S)$ by sending a $(d(S_{\gamma})-1)$-simplex ${\sigma}$ of ${\check{C}}(S_{\gamma})$ to the $(d(S)-1)$-simplex $\xi({\sigma})\cup\{{\gamma}\}$ of $\operatorname{Star}({\gamma})\subset{\check{C}}(S)$. The image of this map is precisely the vertex set of the subgraph $L_{\gamma}$ of ${\check{C}}_{P}(S)$ and it is easy to check that it induces a map between the edges of ${\check{C}}_{P}(S_{\gamma})$ and those of $L_{\gamma}$, from which the conclusion follows. ∎
The procongruence curve complex ${\check{C}}(S)$ can be identified with the simplicial profinite complex whose simplices are the inertia groups $\{{\hat{\mathrm{I}}}_{\sigma}\}_{{\sigma}\in{\check{C}}(S)}\subset{\check{\Gamma}}(S)$ (cf. Remark 4.13 in [6]), so that there is a natural continuous action of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ on ${\check{C}}(S)$:
Lemma 2.9.
For $d(S)\geq 1$, there is a natural continuous monomorphism:
$\check{\Theta}_{\mathbb{I}}\colon\thinspace\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))\hookrightarrow\operatorname{Aut}({\check{C}}(S)).$
Proof.
By Theorem 7.3 in [6], the kernel of the homomorphism $\check{\Theta}_{\mathbb{I}}$ is contained in the subgroup $\operatorname{Hom}(\mathrm{P}{\check{\Gamma}}(S)/Z(\mathrm{P}{\check{\Gamma}}(S)),Z(\mathrm{P}{\check{\Gamma}}(S)))$ of $\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}(S))$ described in Lemma 7.4 and Lemma 3.5 in [6], where we denote by $Z(\mathrm{P}{\check{\Gamma}}(S))$ the center of $\mathrm{P}{\check{\Gamma}}(S)$. From its explicit description, it follows that the intersection of $\operatorname{Hom}(\mathrm{P}{\check{\Gamma}}(S)/Z(\mathrm{P}{\check{\Gamma}}(S)),Z(\mathrm{P}{\check{\Gamma}}(S)))$ with the subgroup $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ inside $\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}(S))$ is trivial, which implies the lemma. ∎
There is a natural action on the link $\operatorname{Lk}({\gamma})\cong{\check{C}}(S_{\gamma})$ of the stabilizer $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))_{{\hat{\mathrm{I}}}_{\gamma}}$ of the inertia group ${\hat{\mathrm{I}}}_{\gamma}$ associated to ${\gamma}$ in $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$. We have:
Lemma 2.10.
After identifying $\operatorname{Lk}({\gamma})$ with ${\check{C}}(S_{\gamma})$, the action of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))_{{\hat{\mathrm{I}}}_{\gamma}}$ on $\operatorname{Lk}({\gamma})$ factors through the natural action of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{\gamma}))$ on ${\check{C}}(S_{\gamma})$. The same statement holds after replacing $\mathrm{P}{\check{\Gamma}}(S)$ with $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ and $\mathrm{P}{\check{\Gamma}}(S_{\gamma})$ with $\mathrm{P}{\check{\Gamma}}^{\pm}(S_{\gamma})$.
Proof.
An automorphism of $\mathrm{P}{\check{\Gamma}}(S)$ which preserves the procyclic subgroup ${\hat{\mathrm{I}}}_{\gamma}$ also preserves its centralizer $Z_{\mathrm{P}{\check{\Gamma}}(S)}({\hat{\mathrm{I}}}_{\gamma})$ in $\mathrm{P}{\check{\Gamma}}(S)$ and, since, for ${\sigma}\in\operatorname{Lk}({\gamma})$, the inertia group ${\hat{\mathrm{I}}}_{\sigma}$ is contained in $Z_{\mathrm{P}{\check{\Gamma}}(S)}({\hat{\mathrm{I}}}_{\gamma})$, the action of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))_{{\hat{\mathrm{I}}}_{\gamma}}$ on $\operatorname{Lk}({\gamma})$ factors through the homomorphism induced by restriction:
$\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))_{{\hat{\mathrm{I}}}_{\gamma}}\to\operatorname{Aut}^{\mathbb{I}}(Z_{\mathrm{P}{\check{\Gamma}}(S)}({\hat{\mathrm{I}}}_{\gamma})).$
By Corollary 4.12 in [6], there is a natural isomorphism $Z_{\mathrm{P}{\check{\Gamma}}(S)}({\hat{\mathrm{I}}}_{\gamma})\cong\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}}$, so that we can identify $\operatorname{Aut}^{\mathbb{I}}(Z_{\mathrm{P}{\check{\Gamma}}(S)}({\hat{\mathrm{I}}}_{\gamma}))$ with the closed subgroup $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}})$ of $\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}})$ consisting of those automorphisms which preserve the set of conjugacy classes of the procyclic subgroups of $\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}}$ generated by profinite Dehn twists.
By Theorem 4.10 in [6], there are the exact sequences:
[…]
where the homomorphism $\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}}\to\{\pm 1\}$ is induced by the action of the stabilizer $\mathrm{P}\Gamma(S)_{\gamma}$ on the orientation of the simple closed curve ${\gamma}$.
Since $\mathrm{P}{\check{\Gamma}}(S)_{\vec{{\gamma}}}$ is the normal subgroup of $\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}}$ topologically generated by Dehn twists, an element of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}})$ preserves the subgroup $\mathrm{P}{\check{\Gamma}}(S)_{\vec{{\gamma}}}$, so that there is a natural homomorphism:
$\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}})\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S)_{\vec{\gamma}}).$
By Theorem 4.14 in [6], the procyclic subgroup ${\hat{\mathrm{I}}}_{\gamma}$ is the center of $\mathrm{P}{\check{\Gamma}}(S)_{\vec{\gamma}}$, hence, it is preserved by every element of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S)_{\vec{\gamma}})$ and there is a natural homomorphism:
$\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S)_{\vec{\gamma}})\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{\gamma})).$
By composing all the above homomorphisms, we get a natural homomorphism:
$\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))_{{\hat{\mathrm{I}}}_{\gamma}}\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{\gamma})).$
The natural isomorphism $\operatorname{Lk}({\gamma})\cong{\check{C}}(S_{\gamma})$ identifies all the inertia subgroups of $\mathrm{P}{\check{\Gamma}}(S)$ contained in $Z_{\mathrm{P}{\check{\Gamma}}(S)}({\hat{\mathrm{I}}}_{\gamma})\cong\mathrm{P}{\check{\Gamma}}(S)_{{\gamma}}$, but which do not contain the Dehn twist $\tau_{\gamma}$, with the inertia subgroups of $\mathrm{P}{\check{\Gamma}}(S_{\gamma})$ in a way which is clearly compatible with the above series of homomorphisms. The first statement of the lemma follows. The second can be proved in a similar way. ∎
The following simple lemma in group theory will also be useful:
Lemma 2.11.
Let $1\to H\to G\to L\to 1$ be a short exact sequence of groups and let $f$ be an automorphism of $H$ such that:
  1. (i)
    the center of $H$ is trivial;
  2. (ii)
    the image $\bar{f}$ of $f$ in $\operatorname{Out}(H)$ normalizes the image of the outer representation $\rho\colon\thinspace L\to\operatorname{Out}(H)$ associated to the given short exact sequence;
  3. (iii)
    the automorphism of $\rho(L)$ induced by the restriction of $\operatorname{inn}\bar{f}$ lifts to an automorphism of $L$.
Then, $f$ extends to an automorphism of $G$.
Proof.
For an element $f\in\operatorname{Aut}(H)$, we denote by $\bar{f}$ its image in $\operatorname{Out}(H)$. Let then $\mathrm{Comp}(L,H)$ be the closed subgroup of $\operatorname{Aut}(L)\times\operatorname{Aut}(H)$ formed by the pairs $(\psi,f)$ such that, for all $\alpha\in L$, there holds (in $\operatorname{Out}(H)$):
$\bar{f}\rho(\alpha)\bar{f}^{-1}=\rho(\psi(\alpha)).$
Since, by hypothesis (i), the center of $H$ is trivial, according to Wells’ exact sequence (cf. Theorem in [25]), there is a canonical isomorphism:
$\operatorname{Aut}(G)_{H}\cong\mathrm{Comp}(L,H),$
where $\operatorname{Aut}(G)_{H}$ is the subgroup of $\operatorname{Aut}(G)$ consisting of those automorphisms which preserve $H$. This isomorphism sends an element $\tilde{f}\in\operatorname{Aut}(G)_{H}$ to the pair $(\psi,f)$, where $\psi\in\operatorname{Aut}(L)$ is the automorphism induced by $\tilde{f}$ passing to the quotient by the normal subgroup $H$ and $f$ is the restriction of $\tilde{f}$ to $H$.
The conclusion follows if we show that an $f\in\operatorname{Aut}(H)$, which satisfies the hypotheses (ii) and (iii) of the lemma, is part of a compatible pair. Since $\operatorname{inn}\bar{f}$ preserves the subgroup $\rho(L)$ and the induced automorphism lifts to $\phi\in\operatorname{Aut}(L)$, it is clear that $(\phi,f)$ is such a compatible pair. ∎
2.6. Connectedness of various curve complexes
Before to proceed with the proof of Proposition 2.1, we need to prove that some curve complexes are connected.
Let $C_{b}(S_{0,n})$ be the curve complex defined as the full subcomplex of the complex of curves $C(S_{0,n})$, for $n\geq 4$, whose vertices consist of isotopy classes of simple closed curves on $S_{0,n}$ which bound a $2$-punctured disc:
Lemma 2.12.
For $n\geq 5$, the simplicial complex $C_{b}(S_{0,n})$ is connected.
Proof.
This can be proved by the same argument which proves the connectivity of the standard curve complex (cf., for instance, the proof of Theorem 4.3 in [9]). We use induction on the geometric intersection number of two simple closed curves $a$ and $b$ which bound a $2$-punctured disc on $S_{0,n}$. When they are disjoint there is nothing to prove. Let us then assume that $a$ and $b$ have geometric intersection $i(a,b)>0$. We claim that there is a simple closed curve $c$ (bounding a $2$-punctured disc) such that $i(a,c)=0$ and $i(c,b)<i(a,b)$.
First, we construct a simple closed curve $c^{\prime}$ (not necessarily bounding a $2$-punctured disc), such that $i(a,c^{\prime})=0$ and $i(c^{\prime},b)<i(a,b)$, by following the oriented path $a$ until it intersects $b$, further following $b$ until it intersects again $a$ for the first time and eventually continuing along the path $a$ in order to close the loop. There are several possibilities depending on the orientations of $a$ and $b$, but, in the end, we find an essential simple closed curve with the required properties.
Now, $c^{\prime}$ bounds a $k$-punctured disk $D$, where $k\geq 2$, which does not contain $a$. If $k=2$, then we take $c=c^{\prime}$ and we are done. Otherwise, we take for $c$ the boundary of a $2$-punctured disc $D^{\prime}$ contained in $D$ such that $c$ is the union of an arc contained in $c^{\prime}$ and an arc which is either part of $b\cap D$ or is disjoint from $b$. In both cases, we have that $i(c,b)\leq i(c^{\prime},b)<i(a,b)$ and, obviously, $i(a,c)=0$. ∎
Let then $C_{0b}(S_{1,n})$ be the full subcomplex of the curve complex $C(S_{1,n})$ whose vertices are isotopy classes of either nonseparating simple closed curves or simple closed curves which bound a $2$-punctured disc on $S_{0,n}$:
Lemma 2.13.
For $n\geq 2$, the simplicial complex $C_{0b}(S_{1,n})$ is connected.
Proof.
We proceed as in the genus zero case. Here, the curve $c$ is either nonseparating or it bounds a $2$-punctured disk but the proof above works without any essential change. ∎
2.7. Proof of Proposition 2.1 for $g(S)=0$
We proceed by induction on $n\geq 5$. The case $n=5$ was proved above and serves as base for the induction. Let us then assume that Proposition 2.1 holds for $S_{0,n-1}$ and let us prove it for $S_{0,n}$.
For $n\geq 5$, every simplex ${\sigma}\in O(S_{0,n})$ contains at least a simple closed curve ${\gamma}$ on $S_{0,n}$ which bounds a $2$-punctured disc. We then have:
Lemma 2.14.
If an automorphism $\phi\in\operatorname{Aut}({\check{C}}_{P}(S_{0,n}))$ preserves the orientation of the profinite Farey subgraph ${\widehat{\operatorname{F}}}_{\sigma}$ of ${\check{C}}_{P}(S_{0,n})$ (cf. Section 2.2), then it preserves the orientations of all profinite subgraphs ${\widehat{\operatorname{F}}}_{{\sigma}^{\prime}}$ such that ${\gamma}\in{\sigma}^{\prime}$, where ${\gamma}\in{\sigma}$ is a simple closed curve which bounds a $2$-punctured disc.
Proof.
By Theorem 6.7 in [6], for all hyperbolic surfaces $S$, there is a natural monomorphism:
$\check{\Theta}_{P}\colon\thinspace\operatorname{Aut}({\check{C}}_{P}(S))\hookrightarrow\operatorname{Aut}({\check{C}}(S)),$
which is induced by the identification of the vertices of ${\check{C}}_{P}(S)$ with the facets of ${\check{C}}(S)$. For $d(S)>1$, one further observes that the continuous action of $\operatorname{Aut}({\check{C}}_{P}(S))$ on the profinite set of Farey subgraphs of ${\check{C}}_{P}(S)$, which are parameterized by the $(d(S)-2)$-simplices of ${\check{C}}(S)$, does indeed induce a continuous action on this profinite set of $(d(S)-2)$-simplices, which is compatible with the action on the facets of ${\check{C}}(S)$ considered above. One then shows that an automorphism of ${\check{C}}(S)$ can be reconstructed from these data.
Thus, after possibly composing the given automorphism $\phi\in\operatorname{Aut}({\check{C}}_{P}(S_{0,n}))$ with an element in the image of $\operatorname{Inn}({\check{\Gamma}}(S_{0,n}))$, we can assume that its image $\check{\Theta}_{P}(\phi)$ in $\operatorname{Aut}({\check{C}}(S_{0,n}))$ preserves the 00-simplex $\{{\gamma}\}\in{\check{C}}(S_{0,n})$ and so $\phi$ preserves the subgraph $L_{\gamma}$ of ${\check{C}}_{P}(S_{0,n})$. Since, by Lemma 2.8, we have that $L_{\gamma}\cong{\check{C}}_{P}(S_{\gamma})$, from the induction hypothesis, it follows that, if the automorphism $\phi$ preserves the orientation of some profinite Farey subgraph of $L_{\gamma}$, then $\phi$ preserves the orientation of all profinite Farey subgraphs ${\widehat{\operatorname{F}}}_{\sigma}$ such that ${\gamma}\in{\sigma}$. ∎
By Lemma 2.12, the curve complex $C_{b}(S_{0,n})$ is connected. Thus, there is a set ${\gamma}_{1},\dots,{\gamma}_{k}$ of simple closed curves on $S_{0,n}$ bounding a $2$-punctured disc such that any two representatives ${\sigma}$ and ${\sigma}^{\prime}$ of the set of orbits $O(S_{0,n})$ are contained in a chain $\operatorname{Lk}({\gamma}_{1}),\dots,\operatorname{Lk}({\gamma}_{k})$ with the property that the intersection $L_{{\gamma}_{i}}\cap L_{{\gamma}_{i+1}}$, for $11\leq i\leq k-$ 1, contains at least a profinite Farey subgraph. Lemma 2.14 and a simple induction then imply that an automorphism $\phi\in\operatorname{Aut}({\check{C}}_{P}(S_{0,n}))$, which preserves the orientation of ${\widehat{\operatorname{F}}}_{\sigma}$, also preserves the orientation of ${\widehat{\operatorname{F}}}_{{\sigma}^{\prime}}$. This completes the proof of Proposition 2.1 for $g(S)=0$.
2.8. Proof of Proposition 2.1 for $g(S)\geq 1$
Let us first consider the case $g(S)=1$. Here, we need to use the curve complex $C_{0b}(S_{1,n})$ instead of the curve complex $C_{b}(S_{0,n})$ and Lemma 2.13 instead of Lemma 2.12. We then proceed by induction on $n\geq 2$. The base of the induction is provided by the case $S=S_{1,2}$ proved above. The rest of the argument proceeds similarly to the case $g(S)=0$ where, besides the induction hypothesis, we also use the case $g(S)=0$ of Proposition 2.1 proved in Section 2.7.
For $g(S)\geq 2$, we proceed by induction on the genus where the base of the induction is the genus $1$ case proved above. The relevant curve complex here is the complex of nonseparating curves $C_{0}(S)$, which, for $g(S)\geq 2$, (cf. Theorem 4.4 in [9]) is connected. The rest of the argument proceeds as in the previous cases.
2.9. A rigidity criterion
From Proposition 8.2 in [6] and Proposition 2.1, it follows that, for $d(S)>1$ and $S\neq S_{1,2}$, there is a natural isomorphism:
$\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\cong\operatorname{Aut}({\check{C}}_{P}(S)).$
For $S=S_{1,2}$, the group $\operatorname{Aut}({\check{C}}_{P}(S_{1,2}))$ must be replaced with the subgroup of those automorphisms preserving the set of nonseparating curves. We will show that this isomorphism implies a characterization of those elements of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ which are induced by an inner automorphism of ${\check{\Gamma}}^{\pm}(S)$. Before we state the result, we need to make a remark.
In the group-theoretic realization of the procongruence curve complex ${\check{C}}(S)$ which we described in Section 2.5, the vertices of the procongruence pants complex ${\check{C}}_{P}(S)$ are identified with the set $\{{\hat{\mathrm{I}}}_{\sigma}\}_{{\sigma}\in{\check{C}}(S)_{d(S)-1}}$ of inertia subgroups of $\mathrm{P}{\check{\Gamma}}(S)$ of maximal rank. The natural faithful continuous action of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ on the curve complex ${\check{C}}(S)$ (cf. Lemma 2.9) then induces a continuous faithful action of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ on the vertex set ${\check{C}}_{P}(S)_{0}$ of the procongruence pants complex. We then have:
Theorem 2.15.
Let $S$ be a connected hyperbolic surface such that $d(S)>1$. An element $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ is in the image of $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ if and only if, for some edge $\{v_{0},v_{1}\}\in{\check{C}}_{P}(S)_{1}$, the set of vertices $\{f(v_{0}),f(v_{1})\}$ is also an edge of ${\check{C}}_{P}(S)$.
2.10. Proof of Theorem 2.15 for $S=S_{0,5}$
Let $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ be an element satisfying the hypotheses of the theorem. Since, for $S=S_{0,5}$, the action of ${\check{\Gamma}}(S)$ on the oriented edges of ${\check{C}}_{P}(S)$ is transitive, after composing with an element in the image of $\operatorname{Inn}({\check{\Gamma}}(S))$, we can assume that $f$ fixes the vertices of the edge $\{v_{0},v_{1}\}$ of ${\check{C}}_{P}(S)$. For the same reason, for any edge $\{\alpha_{0},\alpha_{1}\}$ of ${\check{C}}_{P}(S)$, there is an element $x\in{\check{\Gamma}}(S)$ such that $\alpha_{i}=x\cdot v_{i}\cdot x^{-1}$, for $i=0,1$.
Let us consider the short exact sequence $1\to\mathrm{P}{\check{\Gamma}}(S)\to{\check{\Gamma}}(S)\to\Sigma_{n}\to 1$, where we put $n:=n(S)$. From Corollary C in [11], it follows that the outer representation associated to this short exact sequence identifies $\Sigma_{n}$ with a normal subgroup of $\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$. Since $\mathrm{P}{\check{\Gamma}}(S)$ is center free, from Lemma 2.11, it follows that the given element $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ extends to an automorphism of ${\check{\Gamma}}(S)$, which we also denote by $f$, so that we have:
$f(\alpha_{i})=f(x\cdot v_{i}\cdot x^{-1})=f(x)\cdot v_{i}\cdot f(x)^{-1},\hskip 14.22636pt\mbox{ for }i=0,1,$
and then $\{f(\alpha_{0}),f(\alpha_{1})\}=\operatorname{inn}f(x)(\{v_{0},v_{1}\})\in{\check{C}}_{P}(S)$. Therefore, the continuous action of the automorphism $f$ on the profinite set of vertices of ${\check{C}}_{P}(S)$ extends to a continuous action on the procongruence pants complex ${\check{C}}_{P}(S)$. The conclusion then follows from Proposition 2.1.
2.11. Proof of Theorem 2.15 for $S=S_{1,2}$
To deal with this case, we need first to prove the following lemmas:
Lemma 2.16.
For $S=S_{1,2}$, the action of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$ on ${\check{C}}(S_{1,2})$ preserves topological types. Moreover, there is a natural monomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\hookrightarrow\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$ induced by restriction of automorphisms.
Proof.
With the notation of Section 2.5, for ${\gamma}$ a simple closed curve on $S_{1,2}$, by Corollary 4.12 in [6], there is a natural isomorphism $Z_{\mathrm{P}{\check{\Gamma}}(S_{1,2})}({\hat{\mathrm{I}}}_{\gamma})\cong\mathrm{P}{\check{\Gamma}}(S_{1,2})_{{\gamma}}$ and, by Theorem 4.10 in [6], there are exact sequences:
$1\rightarrow\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}}\to\mathrm{P}{\check{\Gamma}}(S_{1,2})_{{\gamma}}\to\{\pm 1\}\hskip 14.22636pt\mbox{ and }\hskip 14.22636pt1\rightarrow\hat{\mathrm{I}}_{\gamma}\to\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}}\to\mathrm{P}{\check{\Gamma}}((S_{1,2})_{\gamma})\to 1$
After possibly composing with an inner automorphism, we can assume that a given element $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$ is such that ${\gamma}^{\prime}:=f({\gamma})$ also belongs to $C(S_{1,2})_{0}\subset{\check{C}}(S)_{0}$. Since $\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}}$ identifies with the subgroup of the centralizer $Z_{\mathrm{P}{\check{\Gamma}}(S_{1,2})}({\hat{\mathrm{I}}}_{\gamma})$ topologically generated by profinite Dehn twists, we have that $f(\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}})=\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}^{\prime}}$. By Theorem 4.14 in [6], the procyclic subgroup ${\hat{\mathrm{I}}}_{\gamma}$ is the center of $\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}}$, so that $f$ induces an isomorphism $\bar{f}\colon\thinspace\mathrm{P}{\check{\Gamma}}((S_{1,2})_{\gamma})\stackrel{{\scriptstyle\sim}}{{\to}}\mathrm{P}{\check{\Gamma}}((S_{1,2})_{{\gamma}^{\prime}})$.
For ${\gamma}$ separating, we have that $\mathrm{P}{\check{\Gamma}}((S_{1,2})_{\gamma})\cong\widehat{\operatorname{SL}(2,\mathbb{Z})}$ while, for ${\gamma}$ nonseparating, we have that $\mathrm{P}{\check{\Gamma}}((S_{1,2})_{\gamma})$ is a free group in two generators. The latter profinite group is torsion free while the former is not. Thus, ${\gamma}^{\prime}$ has the same topological type of ${\gamma}$.
This proves the first part of the lemma. Let us then observe that, by Lemma 2.5, $\mathrm{P}{\check{\Gamma}}(S_{0,5})$ identifies with the normal subgroup of $\mathrm{P}{\check{\Gamma}}(S_{1,2})$ topologically generated by squares of nonseparating Dehn twists. By the previous part of the proof, elements of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$ preserve this subgroup and so there is a homomorphism as claimed in the lemma. The fact that this is injective follows from the fact that the monomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\hookrightarrow\operatorname{Aut}({\check{C}}(S_{1,2}))$ (cf. Lemma 2.9) factors through it and the monomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\hookrightarrow\operatorname{Aut}({\check{C}}(S_{0,5}))$, via the isomorphism ${\check{C}}(S_{1,2})\cong{\check{C}}(S_{0,5})$. ∎
Since all the groups involved are center free, for $n\geq 4$, there are a series of natural isomorphisms:
$\operatorname{Inn}({\check{\Gamma}}(S_{0,n}))\left/\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))\right.\cong{\check{\Gamma}}(S_{0,n})\left/\mathrm{P}{\check{\Gamma}}(S_{0,n})\right.\cong\Sigma_{n}.$
Let us then denote by $\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$ the centralizer of the image of $\Sigma_{n}$ in $\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$. Since $\Sigma_{n}$ acts transitively on the set of conjugacy classes of procyclic subgroups of $\mathrm{P}{\check{\Gamma}}(S_{0,n})$ generated by Dehn twists about simple closed curves bounding $2$-punctured discs in $S_{0,n}$, it follows that $\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$ preserves such conjugacy classes. Hence, for $n\leq 5$, our definition of this group agrees with the one given in Section 0.1 of [10].
From the isomorphism $\mathrm{P}\Gamma(S_{1,2})\cong{\Gamma}(S_{0,5})_{Q}$ (cf. Section 2.4), it follows that there is a natural isomorphism $\operatorname{Inn}(\mathrm{P}\Gamma(S_{1,2}))\left/\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\right.\cong\Sigma_{4}$. Let us then define $\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$ to be the centralizer of the image of $\Sigma_{4}$ in $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\left/\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\right.$.
For $S$ one of the surfaces considered above, let also $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S))$ be the inverse image of $\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S))$ in $\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}(S))$ (note that $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S))$ is contained in $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$). We then have:
Lemma 2.17.
There are natural isomorphisms:
$\begin{array}[]{l}\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\left/\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\right.\cong\Sigma_{4}\times\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2})),\\ \\ \operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\cong\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2})),\\ \\ \operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\cong\Sigma_{5}\times\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,5})).\end{array}$
Moreover, the natural monomorphism of Lemma 2.16 restricts to a monomorphism:
$\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\hookrightarrow\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,5})).$
Proof.
There are short exact sequences:
$1\to\Sigma_{4}\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\left/\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\right.\to\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\to 1$
and (cf. Corollary C in [11])
$1\to\Sigma_{5}\to\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\to\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\left/\Sigma_{5}\right.\to 1$
Since $\Sigma_{4}$ and $\Sigma_{5}$ are complete groups, the above short exact sequences split and there are natural isomorphisms as stated in the lemma (cf. Theorem 7.15 in [23]). The last statement of the lemma follows from the fact that the centralizer of $\Sigma_{4}$ in $\Sigma_{5}$ is trivial. ∎
Remark 2.18.
By Corollary C in [20] and Main Theorem in [10], there is actually a series of natural isomorphisms $\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\cong\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\cong\widehat{\operatorname{GT}}$, where $\widehat{\operatorname{GT}}$ is the profinite Grothendieck-Teichmüller group, so that the monomorphism of Lemma 2.17 is actually an isomorphism. But we will not need this fact.
Let $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$ be an element such that for some edge $\{v_{0},v_{1}\}\in{\check{C}}_{P}(S_{1,2})_{1}$, the set of vertices $\{f(v_{0}),f(v_{1})\}$ is also an edge of ${\check{C}}_{P}(S_{1,2})$. Let $\tilde{f}$ be the image of $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$ via the monomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\hookrightarrow\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$ of Lemma 2.16.
The element $\tilde{f}$ then acts on the vertex set ${\check{C}}_{P}(S_{0,5})_{0}$ through the element $f$ and the natural continuous ${\check{\Gamma}}(S_{1,2})$-equivariant bijection on vertex sets:
$q\colon\thinspace{\check{C}}_{P}(S_{1,2})_{0}\stackrel{{\scriptstyle\sim}}{{\to}}{\check{C}}_{P}(S_{0,5})_{0},$
so that, for $v\in{\check{C}}_{P}(S_{1,2})_{0}$, there holds $\tilde{f}(q(v))=q(f(v))$. The key lemma is the following:
Lemma 2.19.
The element $\tilde{f}\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$ is such that for some edge $\{w_{0},w_{1}\}\in{\check{C}}_{P}(S_{0,5})_{1}$, the set of vertices $\{\tilde{f}(w_{0}),\tilde{f}(w_{1})\}$ is also an edge of ${\check{C}}_{P}(S_{0,5})$.
Proof.
If the given $\{v_{0},v_{1}\}\in{\check{C}}_{P}(S_{1,2})_{1}$ is an edge contained in a ${\check{\Gamma}}(S_{1,2})$-orbit such that the pair of vertices $\{q(v_{0}),q(v_{1})\}$ is an edge of ${\check{C}}_{P}(S_{0,5})$, then, since $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$, by Lemma 2.16, preserves ${\check{\Gamma}}(S_{1,2})$-orbits, we have that $\{\tilde{f}(q(v_{0})),\tilde{f}(q(v_{1}))\}$ is also an edge of ${\check{C}}_{P}(S_{0,5})$. In this case, we just let $w_{0}:=q(v_{0})$ and $w_{1}:=q(v_{1})$ and we are done. In particular, as it is easy to check, this happens if the common profinite simple closed curve in the intersection $v_{0}\cap v_{1}$ has the topological type of a separating curve on $S_{1,2}$.
Let us then consider the case when the common profinite simple closed curve ${\gamma}$ in $v_{0}\cap v_{1}$ has the topological type of a nonseparating curve on $S_{1,2}$. As usual, it is not restrictive to assume that ${\gamma}\in C(S_{1,2})_{0}\subset{\check{C}}(S_{1,2})_{0}$ and after, possibly, composing the given $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$ with an inner automorphism of $\mathrm{P}\Gamma(S_{1,2})$, we can also assume that $f\in\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$. From Lemma 2.17, it then follows that that the conjugacy class of ${\hat{\mathrm{I}}}_{\gamma}\cap\mathrm{P}{\check{\Gamma}}(S_{0,5})$ in $\mathrm{P}{\check{\Gamma}}(S_{0,5})$ is preserved by $\tilde{f}$. Therefore, after, possibly, composing by an inner automorphism of $\mathrm{P}{\check{\Gamma}}(S_{0,5})$, we can at last assume that $f\in\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}}$.
With the above assumptions, in particular, $f$ preserves the link $\operatorname{Lk}({\gamma})$ of ${\gamma}$ in ${\check{C}}(S_{1,2})$ and then the vertex set of the profinite subgraph $L_{\gamma}$ of ${\check{C}}_{P}(S_{1,2})$. Let us recall (cf. Lemma 2.8) that there are natural $\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\gamma}$-equivariant continuous isomorphisms $\operatorname{Lk}({\gamma})\cong{\check{C}}(S_{1,2}\smallsetminus{\gamma})$ and $L_{\gamma}\cong{\check{C}}_{P}(S_{1,2}\smallsetminus{\gamma})$.
By hypothesis, we have that both $\{v_{0},v_{1}\}$ and $\{f(v_{0}),f(v_{1})\}\in(L_{\gamma})_{1}$. We claim that $f$ preserves all the edge set of $L_{\gamma}$ (and so induces an automorphism of this profinite graph). By Lemma 2.10, $f$ acts on the vertex set $(L_{\gamma})_{0}\cong{\check{C}}_{P}(S_{1,2}\smallsetminus{\gamma})_{0}$ through its image via the natural homomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}}\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma}))$. This is induced by the restriction to the stabilizer $\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{\gamma}}$ followed by the projection to its quotient $\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma})$.
By Lemma 2.17, there is a natural homomorphism $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}}\hookrightarrow\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$, induced by restriction. Since $\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{\gamma}}\cap\mathrm{P}{\check{\Gamma}}(S_{0,5})=\mathrm{P}{\check{\Gamma}}(S_{0,5})_{q({\gamma})}$ and the natural projection $\mathrm{P}{\check{\Gamma}}(S_{0,5})_{q({\gamma})}\to\mathrm{P}{\check{\Gamma}}(S_{0,5}\smallsetminus q({\gamma}))$, after identifying $\mathrm{P}{\check{\Gamma}}(S_{0,5}\smallsetminus q({\gamma}))$ with $\mathrm{P}{\check{\Gamma}}(S_{0,4})$, is just the restriction of a forgetful homomorphism $\mathrm{P}{\check{\Gamma}}(S_{0,5})\to\mathrm{P}{\check{\Gamma}}(S_{0,4})$ to $\mathrm{P}{\check{\Gamma}}(S_{0,5})_{q({\gamma})}$, we see that the restriction of the homomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}}\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma}))$ to the subgroup $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}}$ is equivalent to the restriction of the homomorphism $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\to\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,4}))$, defined in Section 2.2 in [10], associated to the forgetful homomorphism considered above, to the image of $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}}$ in $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$.
In particular, the image of $f$ in $\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma}))$ commutes with the image of $\Sigma_{4}$ via the outer representation associated to the short exact sequence:
$1\to\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma})\to{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma})\to\Sigma_{4}\to 1$
It is then clear that all hypotheses of Lemma 2.11 are satisfied and $f$ extends to an automorphism of ${\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma})$. As in Section 2.10, we then conclude that $f$ induces an automorphism of the pants complex ${\check{C}}_{P}(S_{1,2}\smallsetminus{\gamma})$, as claimed above.
It is now easy to check that, for some edge $\{v_{0}^{\prime},v_{1}^{\prime}\}\in(L_{\gamma})_{1}\cong{\check{C}}_{P}(S_{1,2}\smallsetminus{\gamma})_{1}$, we have that $\{q(v_{0}^{\prime}),q(v_{1}^{\prime})\}\in{\check{C}}_{P}(S_{0,5})_{1}$ and conclude, as we did at the beginning of the proof, letting $w_{0}:=q(v_{0}^{\prime})$ and $w_{1}:=q(v_{1}^{\prime})$. ∎
From Lemma 2.19 and the case $S=S_{0,5}$ of Theorem 2.15 proved above, we conclude that the image $\tilde{f}$ of $f$ in $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$ is in the image of $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S_{0,5}))$. In conclusion, $\tilde{f}$ is an inner automorphism of ${\check{\Gamma}}^{\pm}(S_{0,5})$ which normalizes its subgroup $\mathrm{P}{\check{\Gamma}}(S_{1,2})$.
From Lemma 9.13 in [6], it now follows that the normalizer of $\mathrm{P}{\check{\Gamma}}(S_{1,2})$ in ${\check{\Gamma}}^{\pm}(S_{0,5})$ coincides with the closure in this group of the normalizer of $\mathrm{P}\Gamma(S_{1,2})$ in ${\Gamma}^{\pm}(S_{0,5})$, which, as it easily follows from (ii) of Theorem in [17], is just $\mathrm{P}\Gamma^{\pm}(S_{1,2})$. Since $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S_{1,2}))=\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2}))$, this implies Theorem 2.15 for $S=S_{1,2}$.
2.12. Proof of Theorem 2.15 for $d(S)>2$
We proceed by induction on $d(S)$. Let us then assume that the statement of the lemma holds for all surfaces of modular dimension $<d(S)$. It is clearly not restrictive to assume that the edge $\{v_{0},v_{1}\}$ in the hypothesis of the theorem belongs to $C_{P}(S)\subset{\check{C}}_{P}(S)$.
By Proposition 2.1, it is enough to prove that the action of the given $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ on the set of vertices of the procongruence pants complex ${\check{C}}_{P}(S)$ preserves its set of edges, that is to say, for every edge $\{\alpha_{0},\alpha_{1}\}\in{\check{C}}_{P}(S)_{1}$, there holds $\{f(\alpha_{0}),f(\alpha_{1})\}\in{\check{C}}_{P}(S)_{1}$. By the same argument of the proof of the case $S=S_{0,5}$ of the theorem, it is enough to show that this is the case for a set of representatives of the $\mathrm{P}{\check{\Gamma}}(S)$-orbits in ${\check{C}}_{P}(S)_{1}$. In particular, we can assume that $\{\alpha_{0},\alpha_{1}\}\in C_{P}(S)_{1}\subset{\check{C}}_{P}(S)_{1}$ as well.
For $d(S)>2$, the complexes $C_{b}(S)$ (for $g(S)=0$), $C_{0b}(S)$ (for $g(S)=1$) and $C_{0}(S)$ (for $g(S)>1$) are connected. This implies that there is a set ${\gamma}_{1},\dots,{\gamma}_{k}$ of simple closed curves on $S$, where ${\gamma}_{i}$, for $i=1,\dots,k$, is either nonseparating or bounds a $2$-punctured disc, such that the edge $\{v_{0},v_{1}\}$ is contained in $L_{{\gamma}_{1}}$, the edge $\{\alpha_{0},\alpha_{1}\}$ is contained in $L_{{\gamma}_{k}}$ and the intersection $L_{{\gamma}_{i}}\cap L_{{\gamma}_{i+1}}$, for $11\leq i\leq k-$ 1, contains at least an edge of ${\check{C}}_{P}(S)$.
The conclusion then follows from a simple induction and the following lemma:
Lemma 2.20.
If an automorphism $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ sends an edge of $L_{{\gamma}_{i}}$ to an edge of ${\check{C}}_{P}(S)$, then it sends every edge of $L_{{\gamma}_{i}}$ to an edge of ${\check{C}}_{P}(S)$, for $i=1,\dots,k$.
Proof.
After composing $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ with an element in the image of $\operatorname{Inn}({\check{\Gamma}}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$, we can assume that $f$ preserves the procyclic subgroup ${\hat{\mathrm{I}}}_{{\gamma}_{i}}$ and then acts on the vertex set of the subgraph $L_{{\gamma}_{i}}$, which identifies with the set of $(d(S)-1)$-simplices of the star of ${\gamma}_{i}$ in ${\check{C}}(S)$.
By Lemma 2.8, the profinite subgraph $L_{{\gamma}_{i}}$ is naturally isomorphic to ${\check{C}}_{P}(S_{{\gamma}_{i}})$ and, by Lemma 2.10, this natural isomorphism induces an action of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))_{{\hat{\mathrm{I}}}_{{\gamma}_{i}}}$ on the vertex set of ${\check{C}}_{P}(S_{{\gamma}_{i}})$ which factors through an element of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{{\gamma}_{i}}))$. The induction hypothesis then implies that $f$ preserves the edge set of $L_{{\gamma}_{i}}$, for $i=1,\dots,k$. ∎
3. Antiholomorphic involutions
3.1. Centralizers of antiholomorphic involutions
An antiholomorphic involution $\iota\in{\Gamma}^{\pm}(S)$ is an element of order $2$ (an involution) which reverses the orientation of $S$. Any such element can be realized as the antiholomorphic involution associated to a real Riemann surface homeomorphic to $S$.
The centralizer of $\iota$ in ${\Gamma}^{\pm}(S)$ has a simple description. Let $S_{\iota}:=S/\langle\iota\rangle$ be the quotient surface. Let $\operatorname{Fix}(\iota)$ be the fixed point set of $\iota$. Then, $\operatorname{Fix}(\iota)$ is the union of a (possibly empty) set of disjoint simple closed curves on $S$ and the quotient surface $S_{\iota}$ is orientable if and only if $S\smallsetminus\operatorname{Fix}(\iota)$ is not connected. Moreover, if $\operatorname{Fix}(\iota)\neq\emptyset$, then $S_{\iota}$ is a surface with boundary $\partial S_{\iota}$, which coincides with the image of $\operatorname{Fix}(\iota)$ in $S_{\iota}$ (cf. Proposition 1.2 in [24]). Let us denote by $\operatorname{Map}(S_{\iota})$ the group of isotopy classes of self-diffeomorphisms of the (possibly non-orientable) surface $S_{\iota}$. We then have:
Proposition 3.1.
The centralizer $Z_{{\Gamma}^{\pm}(S)}(\iota)$ of $\iota$ in ${\Gamma}^{\pm}(S)$ is described by the short exact sequence:
$1\to\langle\iota\rangle\to Z_{{\Gamma}^{\pm}(S)}(\iota)\to\operatorname{Map}(S_{\iota})\to 1$
Proof.
If $\operatorname{Fix}(\iota)=\emptyset$, it is enough to observe that the orientation cover $S\to S_{\iota}$ is canonical. This implies that any self-homeomorphism of $S_{\iota}$ lifts to $S$ and so the proposition follows in this case.
Let us then assume that $\operatorname{Fix}(\iota)\neq\emptyset$ and $S\smallsetminus\operatorname{Fix}(\iota)$ is connected. The surface $S\smallsetminus\operatorname{Fix}(\iota)$ identifies with the orientation cover of $S_{\iota}\smallsetminus\partial S_{\iota}$, so that every self-homeomorphism of $S_{\iota}\smallsetminus\partial S_{\iota}$ lifts to $S\smallsetminus\operatorname{Fix}(\iota)$. Since every self-homeomorphism of $S$ which commutes with $\iota$ preserves the fixed point set $\operatorname{Fix}(\iota)$, the conclusion follows.
Let us then consider the case when $S\smallsetminus\operatorname{Fix}(\iota)$ is not connected. In this case, $S\smallsetminus\operatorname{Fix}(\iota)$ has two connected components $S^{\prime}$ and $S^{\prime\prime}$ such that their closures $\overline{S}^{\prime}$ and $\overline{S}^{\prime\prime}$ in $S$ both identify with the quotient surface $S_{\iota}$. This implies that a self-homeomorphism of $S_{\iota}$ lifts to a pair of self-homeomorphisms of $\overline{S}^{\prime}$ and $\overline{S}^{\prime\prime}$ which are compatible on the boundary and can then be glued to a self-homeomorphism of $S$. ∎
3.2. The fixed point set of an antiholomorphic involution in the augmented Teichmüller space
Let ${\mathcal{T}}(S)$ be the Teichmüller space associated to the surface $S$ endowed with the Weil-Petersson metric. From Lemma 3.5 in [22], it follows that the fixed point set ${\mathcal{T}}(S)^{\iota}$ of an antiholomorphic involution $\iota\in{\Gamma}^{\pm}(S)$ is a nonempty and connected real submanifold of ${\mathcal{T}}(S)$ of (real) dimension $d(S)$ (cf. Corollary 3.8 in [22]).
Since the augmented Teichmüller space $\overline{{\mathcal{T}}}(S)$ is the completion of the Teichmüller space ${\mathcal{T}}(S)$ with respect to the Weil-Petersson metric (cf. Section 2.1), from Theorem 5 in [26], it follows that the fixed point set $\overline{{\mathcal{T}}}(S)^{\iota}$ coincides with the closure of the fixed point set ${\mathcal{T}}(S)^{\iota}$ in $\overline{{\mathcal{T}}}(S)$.
Theorem 5 in [26] and Lemma 3.5 in [22] then imply that, for ${\sigma}\in C(S)_{k}$, the fixed point set $\partial\overline{{\mathcal{T}}}(S)_{\sigma}^{\iota}$ of the corresponding closed stratum of $\partial\overline{{\mathcal{T}}}(S)$ is nonempty if and only if ${\sigma}\in C(S)_{k}^{\iota}$ and that, for all ${\sigma}\in C(S)_{k}^{\iota}$, there holds $\partial\overline{{\mathcal{T}}}(S)_{\sigma}^{\iota}\cong\overline{{\mathcal{T}}}(S\smallsetminus{\sigma})^{\iota}$. We sum up the above discussion in the following proposition:
Proposition 3.2.
The closed strata of codimension $k+1$ in the boundary of the fixed point locus $\overline{{\mathcal{T}}}(S)^{\iota}$ are parameterized by the fixed point set $C(S)_{k}^{\iota}$, for $k\geq 0$.
Remark 3.3.
Note that, for all $k\geq 0$, the action of the centralizer $Z_{{\Gamma}^{\pm}(S)}(\iota)$ on $C(S)_{k}$ preserves the fixed point set $C(S)_{k}^{\iota}$ and acts on the latter with a finite number of orbits.
3.3. Antiholomorphic involutions of the profinite mapping class group
Let us now assume that $g(S)\leq 2$. By the congruence subgroup property in genus $\leq 2$, we then have ${\check{\Gamma}}(S)\cong{\widehat{\Gamma}}(S)$ and so ${\check{\Gamma}}^{\pm}(S)\cong{\widehat{\Gamma}}^{\pm}(S)$. The augmentation map ${\Gamma}^{\pm}(S)\to\mathbb{Z}/2$ induces an augmentation map ${\widehat{\Gamma}}^{\pm}(S)\to\mathbb{Z}/2$ and we define an antiholomorphic involution of ${\widehat{\Gamma}}^{\pm}(S)$ to be an element of order $2$ whose image by the augmentation map is nontrivial.
It is well known that ${\Gamma}(S)$ is a good group for $g(S)\leq 2$. The same then holds for ${\Gamma}^{\pm}(S)$ and, from Corollary 3.5 in [21], it follows that the natural inclusion ${\Gamma}^{\pm}(S)<{\widehat{\Gamma}}^{\pm}(S)$ induces a bijection between the conjugacy classes of elements of prime order in ${\Gamma}^{\pm}(S)$ and in ${\widehat{\Gamma}}^{\pm}(S)$. In particular, we have:
Proposition 3.4.
For $g(S)\leq 2$, every antiholomorphic involution of ${\widehat{\Gamma}}^{\pm}(S)$ is conjugated to an antiholomorphic involution inside ${\Gamma}^{\pm}(S)$.
3.4. The fixed point set of an antiholomorphic involution in the procongruence moduli stack
A description of the fixed point locus $\overline{{\mathbb{M}}}(S)^{\iota}$ for the action of $\iota$ on $\overline{{\mathbb{M}}}(S)$ for $g(S)\leq 2$ is implicit in the proof of Corollary 3.15 in [22] where it is also proved that the centralizer $Z_{{\widehat{\Gamma}}^{\pm}(S)}(\iota)$ coincides with the closure of $Z_{{\Gamma}^{\pm}(S)}(\iota)$ in ${\widehat{\Gamma}}^{\pm}(S)$. Here, we prefer to give an alternative proof of these statements since the argument can be substantially simplified using Corollary 3.5 in [21]:
Proposition 3.5.
For $g(S)\leq 2$, the fixed point set locus $\overline{{\mathbb{M}}}(S)^{\iota}$ of an antiholomorphic involution $\iota\in{\Gamma}^{\pm}(S)<{\widehat{\Gamma}}^{\pm}(S)$ is smooth, irreducible and contains ${{\mathbb{M}}}(S)^{\iota}$ as an open dense subspace. The centralizer $Z_{{\widehat{\Gamma}}^{\pm}(S)}(\iota)$ coincides with the closure of $Z_{{\Gamma}^{\pm}(S)}(\iota)$ in ${\widehat{\Gamma}}^{\pm}(S)$.
Proof.
For every torsion free characteristic level ${\Gamma}^{\lambda}$ of ${\Gamma}(S)$, let $({\mathcal{M}}(S)^{\lambda},\iota)$ be the real complex manifold with equivariant fundamental group (cf. Section 3 in [12]) isomorphic to ${\Gamma}^{\lambda}\cdot\langle\iota\rangle$. By Theorem 3.6 in [22], there is a natural bijection between the set of connected components of the real locus of $({\mathcal{M}}(S)^{\lambda},\iota)$ and conjugacy classes of involutions in ${\Gamma}^{\lambda}\cdot\langle\iota\rangle$. For $g(S)\leq 2$, the latter is a good group and then, by Corollary 3.5 in [21], we conclude that there is a bijective correspondence between the set of connected components of the real locus of $({\mathcal{M}}(S)^{\lambda},\iota)$ and conjugacy classes of involutions in ${\widehat{\Gamma}}^{\lambda}\cdot\langle\iota\rangle$. Passing to the inverse limit over all such level structures, since $\cap_{{\lambda}\in\Lambda}{\widehat{\Gamma}}^{\lambda}\cdot\langle\iota\rangle=\langle\iota\rangle$, it follows that the fixed point set locus ${\mathbb{M}}(S)^{\iota}$ is connected, smooth (and so irreducible) and that the centralizer $Z_{{\widehat{\Gamma}}^{\pm}(S)}(\iota)$ identifies with the covering transformation group of the cover from ${\mathbb{M}}(S)^{\iota}$ to the connected component of ${\mathcal{M}}(S)_{\mathbb{R}}$ associated to the conjugacy class of $\iota$ in ${\widehat{\Gamma}}^{\pm}(S)$.
The statements in the lemma about the fixed point set locus $\overline{{\mathbb{M}}}(S)^{\iota}$ then follow from the fact that this locus is the closure of ${\mathbb{M}}(S)^{\iota}$ in $\overline{{\mathbb{M}}}(S)$. ∎
The closed boundary strata of $\overline{{\mathbb{M}}}(S)$ are parameterized by the simplices of the procongruence curve complex ${\check{C}}(S)$ and, for ${\sigma}\in C(S)\subset{\check{C}}(S)$, there holds $\partial\overline{{\mathbb{M}}}(S)_{\sigma}\cong\overline{{\mathbb{M}}}(S\smallsetminus{\sigma})$, where we denote by $\partial\overline{{\mathbb{M}}}(S)_{\sigma}$ the closed stratum of $\partial\overline{{\mathbb{M}}}(S)=\overline{{\mathbb{M}}}(S)\smallsetminus{\mathbb{M}}(S)$ parameterized by a simplex ${\sigma}\in{\check{C}}(S)$. From Proposition 3.2 and Proposition 3.5, it then follows:
Proposition 3.6.
For $g(S)\leq 2$, the closed strata of the boundary of the fixed point locus $\overline{{\mathbb{M}}}(S)^{\iota}$ of codimension $k+1$ are parameterized by the fixed point set ${\check{C}}(S)_{k}^{\iota}$, for all $k\geq 0$. Moreover, ${\check{C}}(S)_{k}^{\iota}$ is the closure of the fixed point set $C(S)_{k}^{\iota}$ in the profinite set ${\check{C}}(S)_{k}$.
Proof.
From the proof of Proposition 3.5, it follows that the fixed point locus ${{\mathbb{M}}}(S)^{\iota}$ is the closure of the image of the fixed point locus ${\mathcal{T}}(S)^{\iota}$ in ${{\mathbb{M}}}(S)$ and that the fixed point locus $\overline{{\mathbb{M}}}(S)^{\iota}$ is the closure of the image of the fixed point locus $\overline{{\mathcal{T}}}(S)^{\iota}$ in $\overline{{\mathbb{M}}}(S)$. In particular, the boundary $\partial\overline{{\mathbb{M}}}(S)^{\iota}=\overline{{\mathbb{M}}}(S)^{\iota}\smallsetminus{{\mathbb{M}}}(S)^{\iota}$ is the closure of the boundary $\partial\overline{{\mathcal{T}}}(S)^{\iota}$ in the boundary $\partial\overline{{\mathbb{M}}}(S)=\overline{{\mathbb{M}}}(S)\smallsetminus{{\mathbb{M}}}(S)$. By Proposition 3.2, both claims of the proposition then follow. ∎
4. Proof of Theorem 1.1
4.1. Preliminary lemmas
Before we proceed to the proof of Theorem 1.1, we need to prove first a series of lemmas. A subgroup of ${\check{\Gamma}}^{\pm}(S)$ is $\mathbb{I}$-characteristic if it is preserved by all elements of $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$.
Lemma 4.1.
For all $g(S)\geq 0$, $\mathrm{P}{\check{\Gamma}}(S)$ is an $\mathbb{I}$-characteristic subgroup of $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ and $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ is an $\mathbb{I}$-characteristic subgroup of ${\check{\Gamma}}^{\pm}(S)$.
Proof.
The pure mapping class group $\mathrm{P}\Gamma(S)$ is the subgroup of the mapping class group ${\Gamma}(S)$ generated by Dehn twists. Hence, $\mathrm{P}{\check{\Gamma}}(S)$ is an $\mathbb{I}$-characteristic subgroup of both ${\check{\Gamma}}(S)$ and $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$. In order to prove that $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ is an $\mathbb{I}$-characteristic subgroup of ${\check{\Gamma}}^{\pm}(S)$, it is enough to show that ${\check{\Gamma}}(S)$ is an $\mathbb{I}$-characteristic subgroup of ${\check{\Gamma}}^{\pm}(S)$.
It is then enough to show that $\operatorname{Inn}({\check{\Gamma}}(S))$ is a normal subgroup of $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$. This will follow if we show that, for any essential simple closed curve ${\gamma}$ on $S$, there is a character:
$\chi_{\gamma}\colon\thinspace\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to{\widehat{\mathbb{Z}}}^{\ast},$
whose kernel contains $\operatorname{Inn}({\check{\Gamma}}(S))$ but not $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))$.
There is a natural representation $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}({\check{C}}(S))$, which, by Theorem 5.5 in [6], preserves topological types. Therefore, for any $f\in\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$, there is a $x\in{\check{\Gamma}}(S)$ such that $\operatorname{inn}x\circ f$ preserves the procyclic subgroup ${\hat{\mathrm{I}}}_{\gamma}$ generated by the Dehn twist $\tau_{\gamma}$. If $y\in{\check{\Gamma}}(S)$ is another such element, then $xy^{-1}$ fixes ${\gamma}$ and there holds $\operatorname{inn}(xy^{-1})(\tau_{\gamma})=\tau_{\gamma}$, so that $\operatorname{inn}(xy^{-1})$ acts trivially on ${\hat{\mathrm{I}}}_{\gamma}$.
Therefore, assigning to $f\in\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$ the automorphism induced by $\operatorname{inn}x\circ f$ on the subgroup ${\hat{\mathrm{I}}}_{\gamma}$, defines a representation $\chi_{\gamma}\colon\thinspace\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}({\hat{\mathrm{I}}}_{\gamma})$ which contains $\operatorname{Inn}({\check{\Gamma}}(S))$ in its kernel. On the other hand, for an element $x$ of ${\Gamma}^{\pm}(S)$ which is not orientation preserving, there holds $\chi_{\gamma}(\operatorname{inn}x)(\tau_{\gamma})=\tau_{\gamma}^{-1}$, which implies $\ker\chi_{\gamma}=\operatorname{Inn}({\check{\Gamma}}(S))$. ∎
From Lemma 4.1, it follows that the elements of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ are compatible with the augmentation map $\mathrm{P}{\check{\Gamma}}^{\pm}(S)\to\mathbb{Z}/2$ and, in particular, preserve the set of antiholomorphic involutions. Moreover, by Proposition 3.4, for $g(S)\leq 2$, the sets of conjugacy classes of antiholomorphic involutions in $\mathrm{P}\Gamma^{\pm}(S)$ and $\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)$ can be identified. In particular, we have:
Lemma 4.2.
For $g(S)=0$, there is only one ${\widehat{\Gamma}}(S)^{\pm}$-conjugacy class of antiholomorphic involutions in $\mathrm{P}{\widehat{\Gamma}}(S)^{\pm}$ on which $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\widehat{\Gamma}}(S)^{\pm})$ acts naturally.
Proof.
Antiholomorphic involutions in $\mathrm{P}\Gamma(S)^{\pm}$ do not swap the punctures of $S$. Therefore, all the punctures lie in $\operatorname{Fix}(\iota)$, which, in particular, is not empty. Since $g(S)=0$, $\operatorname{Fix}(\iota)$ is also connected and separating. Hence, an antiholomorphic involution in $\mathrm{P}\Gamma(S)^{\pm}$ is determined by the cyclic order of the punctures on $\operatorname{Fix}(\iota)$, so that two of them are conjugated by an element of ${\Gamma}(S)^{\pm}$. The conclusion then follows from Proposition 3.4. ∎
Lemma 4.3.
For $g(S)=0$, the fixed point set $C(S)_{0}^{\iota}$ of an antiholomorphic involution $\iota\in\mathrm{P}\Gamma(S)^{\pm}$ is finite and consists of isotopy classes of simple closed curves on $S$ which have between them geometric intersection either 00 or $2$.
Proof.
The complement $S\smallsetminus\operatorname{Fix}(\iota)$ is the disjoint union of two unpunctured discs Let then $D^{\prime}$ and $D^{\prime\prime}$ be closed subdiscs of $S$ such that $D^{\prime}\cap D^{\prime\prime}=\operatorname{Fix}(\iota)$. For a representative $\alpha$ of $\{[\alpha]\}\in C(S)_{0}^{\iota}$ such that $\alpha=\iota(\alpha)$, the intersection $\alpha\cap D^{\prime}$ is a disjoint union of arcs with boundary in $\operatorname{Fix}(\iota)$. If $\alpha^{\prime}$ is one of such arcs, then $\alpha^{\prime}\cup\iota(\alpha^{\prime})\subset\alpha$ is a simple closed curve (so that $\alpha^{\prime}\cup\iota(\alpha^{\prime})=\alpha$), which crosses transversally $\operatorname{Fix}(\iota)$ in the two boundary points of $\alpha^{\prime}$. The isotopy class of $\alpha$ is then determined by the partition which $\alpha^{\prime}$ induces on the set of punctures of $S$ (which all lie on $\operatorname{Fix}(\iota)$). This implies the first statement of the proposition.
For $\{[\alpha]\}\neq\{[\beta]\}\in C(S)_{0}^{\iota}$ such that $\beta=\iota(\beta)$, the simple closed curve $\beta$ is the union of the two arcs $\beta^{\prime}=\beta\cap D^{\prime}$ and $\beta^{\prime\prime}=\beta\cap D^{\prime\prime}$ and it is clear that the arcs $\alpha^{\prime}$ and $\beta^{\prime}$ have geometric intersection either 00 or $1$. The second statement of the proposition then follows as well. ∎
By Lemma 4.1, there are natural homomorphisms $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ and $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$. We have:
Lemma 4.4.
The homomorphisms:
  1. (i)
    $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ and
  2. (ii)
    $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$,
induced by restriction of automorphisms, are injective.
Proof.
(i): The statement is trivial for $n(S)\leq 1$ so that we can assume $n(S)>1$. For $S=S_{1,2}$, we have that ${\check{\Gamma}}^{\pm}(S)=\mathrm{P}{\check{\Gamma}}^{\pm}(S)\times\langle{\upsilon}\rangle$, where ${\upsilon}$ is the hyperelliptic involution of ${\Gamma}(S)$. This identity implies the stronger statement that $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$.
Thus, we can assume that the center $Z({\check{\Gamma}}^{\pm}(S))$ of ${\check{\Gamma}}^{\pm}(S)$ is trivial. Since the restriction of the homomorphism $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ to $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))$ is injective, the conclusion then follows from Lemma 3.3 in [6].
(ii): Wells’ exact sequence (cf. Theorem in [25]), applied to the short exact sequence $1\to\mathrm{P}{\check{\Gamma}}(S)\to\mathrm{P}{\check{\Gamma}}^{\pm}(S)\to\mathbb{Z}/2\to 1$, implies that the kernel of the given homomorphism is contained in the subgroup of $\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))_{\mathrm{P}{\check{\Gamma}}(S)}$ determined by the group of homomorphisms $\operatorname{Hom}(\mathbb{Z}/2,Z(\mathrm{P}{\check{\Gamma}}(S)))$. From the explicit description of this subgroup, it follows that, even in the cases when $Z(\mathrm{P}{\check{\Gamma}}(S))$ is not trivial, the intersection of the image of $\operatorname{Hom}(\mathbb{Z}/2,Z(\mathrm{P}{\check{\Gamma}}(S)))$ in $\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))_{\mathrm{P}{\check{\Gamma}}(S)}$ with $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ is trivial, which implies item (ii) of the lemma. ∎
By Lemma 4.4, we then have:
Lemma 4.5.
There is a chain of natural inclusions:
$\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\subseteq\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\subseteq\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))\subseteq\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S)).$
4.2. Proof of Theorem 1.1
By Lemma 4.5, in order to prove Theorem 1.1, it is enough to show that the images of $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))$ and $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ in $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ coincide. The idea is to use Theorem 2.15. We need to consider separately three different cases.
4.3. Proof of Theorem 1.1 for $g(S)=0$
By Lemma 4.2, after composing with an inner automorphism of ${\widehat{\Gamma}}(S)^{\pm}$, we can assume that a given element $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\widehat{\Gamma}}(S)^{\pm})$ preserves a fixed antiholomorphic involution $\iota\in\mathrm{P}\Gamma(S)^{\pm}$. The latter has for fixed point set in the surface $S$ a separating simple closed curve containing all punctures of $S$.
By Lemma 4.3, the fixed point set $C(S)^{\iota}_{0}$, for the action of $\iota$ on $C(S)_{0}$, is finite and then, by Proposition 3.6, identifies with the fixed point set ${\check{C}}(S)^{\iota}_{0}$, for the action of $\iota$ on ${\check{C}}(S)_{0}$.
Let $\{\alpha,\beta\}$ be a pair of $\iota$-invariant simple closed curves on $S$ which intersect precisely in two points. An element $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S))$, such that $f(\iota)=\iota$, then preserves $C(S)^{\iota}_{0}={\check{C}}(S)^{\iota}_{0}$ and sends the pair $\{[\alpha],[\beta]\}$ of 00-simplices in $C(S)^{\iota}_{0}$ to the pair $\{f([\alpha]),f([\beta])\}$ also contained in $C(S)^{\iota}_{0}$. Since $f([\alpha])$ and $f([\beta])$ cannot have trivial geometric intersection, otherwise $f$ would not preserve the simplicial structure of ${\check{C}}(S)$, from Lemma 4.3, it follows that $f([\alpha])$ and $f([\beta])$ have geometric intersection $2$.
We can then complete the two 00-simplices $\{[\alpha]\},\{[\beta]\}\in C(S)^{\iota}_{0}$ to two $(n(S)-4)$-simplices $v_{\alpha},v_{\beta}$ of $C(S)$ whose sets of vertices coincide except for the elements $[\alpha]\in v_{\alpha}$ and $[\beta]\in v_{\beta}$. In this way, we have defined an edge $\{v_{\alpha},v_{\beta}\}$ of the pants complex $C_{P}(S)\subset{\check{C}}_{P}(S)$ with the property that $\{f(v_{\alpha}),f(v_{\beta})\}$ is also an edge of ${\check{C}}_{P}(S)$. By Theorem 2.15, we conclude that $f\in\operatorname{Inn}({\widehat{\Gamma}}^{\pm}(S))$.
4.4. Proof of Theorem 1.1 for $S=S_{1,2}$
By Lemma 2.16 and (ii) of Lemma 4.4, there is a natural monomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2}))\hookrightarrow\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{0,5}))$ induced by restriction of automorphisms. Hence, by the case $S=S_{0,5}$ treated above, we have that $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2}))\subseteq\operatorname{Inn}({\widehat{\Gamma}}^{\pm}(S_{0,5}))$. This implies this case of the theorem, since, as in the proof of the case $S=S_{1,2}$ of Theorem 2.15, by Lemma 9.13 of [6], the inner automorphisms of ${\widehat{\Gamma}}^{\pm}(S_{0,5})$ which normalize its subgroup $\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2})$ belong to $\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2}))$.
4.5. Proof of Theorem 1.1, for $g(S)\geq 1$ and $d(S)>2$
The hypotheses implies that, for a nonseparating simple closed curve ${\gamma}$ on $S$, we have $d(S\smallsetminus{\gamma})>1$. We then proceed by induction on the genus of $S$, where the base for the induction is provided by Section 4.3.
Given an element $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S))$, let us consider its action on ${\check{C}}(S)$. By Theorem 5.5 in [6], after possibly composing with an inner automorphism of $\mathrm{P}{\widehat{\Gamma}}(S)$, we may assume that $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))_{{\hat{\mathrm{I}}}_{\gamma}}$, where ${\hat{\mathrm{I}}}_{\gamma}$ is the procyclic subgroup associated to some nonseparating simple closed curve ${\gamma}$ on $S$. In particular, the automorphism $f$ preserves the star $\operatorname{Star}({\gamma})$ and, by Lemma 2.10, acts on the link $\operatorname{Lk}({\gamma})\cong{\check{C}}(S_{\gamma})$ through its image in the group $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{\gamma})$. Thus, $f$ preserves the vertex set of the subgraph $L_{\gamma}$ of ${\check{C}}_{P}(S)$, which identifies with profinite set of $(d(S)-1)$-simplices of $\operatorname{Star}({\gamma})$, and acts continuously on this set through an element of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{\gamma})$. Since, by Lemma 2.8, $L_{\gamma}\cong{\check{C}}_{P}(S_{\gamma})$, the induction hypothesis implies that $f$ also preserves the edge set of $L_{\gamma}$. By Theorem 2.15, we then conclude that $f\in\operatorname{Inn}({\widehat{\Gamma}}^{\pm}(S))$.
5. Proof of Theorem 1.3
By Theorem 1.1, in order to prove Theorem 1.3, we have to show that, for $g(S)=0$, there holds $\operatorname{Aut}({\widehat{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}({\widehat{\Gamma}}^{\pm}(S))$ and $\operatorname{Aut}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S))$. For this, we need to establish first that $\mathrm{P}{\widehat{\Gamma}}(S)$ is a characteristic subgroup of both $\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)$ and ${\widehat{\Gamma}}^{\pm}(S)$. We have:
Lemma 5.1.
For $n\geq 5$, there are unique surjective homomorphisms ${\Gamma}(S_{0,n})\to\Sigma_{n}$ and ${\Gamma}^{\pm}(S_{0,n})\to\Sigma_{n}$, up to automorphisms of $\Sigma_{n}$.
Proof.
Let $G$ be a group which acts transitively on a set of $n$ letters and is generated by $n-1$ elements satisfying the standard braid relations. By a classical result of Artin (cf. Theorem 3 in [2] and its proof), for $n\geq 4$, there is a unique epimorphism $G\to\Sigma_{n}$, up to automorphisms of $\Sigma_{n}$. This result applies, in particular, to the mapping class group ${\Gamma}(S_{0,n})$, from which, the first statement of the lemma follows.
By Lemma 6 in [2], a group $G$ as above does not admit an epimorphism to the alternating group $A_{n}$, for $n\geq 5$. This implies that there is no epimorphism ${\Gamma}(S_{0,n})\to A_{n}$, for $n\geq 5$. Hence, an epimorphism ${\Gamma}^{\pm}(S_{0,n})\to\Sigma_{n}$ restricts to an epimorphism ${\Gamma}(S_{0,n})\to\Sigma_{n}$. From the first part of the proof, it then follows that the kernel of the epimorphism ${\Gamma}^{\pm}(S_{0,n})\to\Sigma_{n}$ contains the pure mapping class group $\mathrm{P}\Gamma(S_{0,n})$.
We now observe that the quotient ${\Gamma}^{\pm}(S_{0,n})/\mathrm{P}\Gamma(S_{0,n})$ is isomorphic to the direct product $\Sigma_{n}\times\mathbb{Z}/2$ which, for $n\geq 3$, admits a unique epimorphism onto $\Sigma_{n}$, up to automorphisms of the latter group. This proves the second claim of the lemma as well. ∎
We also have the following group-theoretic lemma (cf.  Lemma 2.3 in [20]):
Lemma 5.2.
Let $G$ be a finitely generated group and $V$ a finite index normal subgroup with the property that all epimorphisms from $G$ to the quotient group $G/V$ have the same kernel $V$. Then, the closure $\widehat{V}$ of $V$ in the profinite completion $\widehat{G}$ of $G$ is an open characteristic subgroup.
Proof.
Since $G$ is finitely generated, by a classical result of Nikolov and Segal, any epimorphism $\widehat{G}\to G/V$ is continuous and so restricts to an epimorphism $G\to G/V$, which, by our hypothesis, has kernel $V$. Hence, all epimorphisms from $\widehat{G}$ to $G/V$ have the same kernel $\widehat{V}$, which shows that $\widehat{V}$ is indeed a characteristic subgroup of $\widehat{G}$. ∎
We then have the following refinement of Proposition 4.1 (ii) in [20]:
Proposition 5.3.
For $g(S)=0$, the profinite pure mapping class group $\mathrm{P}{\widehat{\Gamma}}(S)$ is a characteristic subgroup of ${\widehat{\Gamma}}(S)$, $\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)$ and ${\widehat{\Gamma}}^{\pm}(S)$.
Proof.
The fact that $\mathrm{P}{\widehat{\Gamma}}(S)$ is characteristic in $\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)$ follows from the fact that $\mathrm{P}{\widehat{\Gamma}}(S)$ is the only torsion free index $2$ subgroup of $\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)$. For $n(S)\geq 5$, the other assertions are immediate consequences of Lemma 5.1 and Lemma 5.2. For $n(S)=4$, we just observe that $\mathrm{P}{\widehat{\Gamma}}(S)$ is the maximal normal free subgroup contained in all the groups in the statement of the proposition. ∎
We can now prove the following lemma, which, as observed above, implies Theorem 1.3:
Lemma 5.4.
For $g(S)=0$ and $n(S)\geq 5$, there holds $\operatorname{Aut}({\widehat{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}({\widehat{\Gamma}}^{\pm}(S))$ and $\operatorname{Aut}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S))$.
Proof.
By Proposition 5.3, there are natural homomorphisms $\operatorname{Aut}({\widehat{\Gamma}}^{\pm}(S))\to\operatorname{Aut}(\mathrm{P}{\widehat{\Gamma}}(S))$ and $\operatorname{Aut}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S))\to\operatorname{Aut}(\mathrm{P}{\widehat{\Gamma}}(S))$ induced by restriction of automorphisms. The conclusion then follows from the identity $\operatorname{Aut}(\mathrm{P}{\widehat{\Gamma}}(S))=\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\widehat{\Gamma}}(S))$ (cf. Lemma 3.13 in [7] and Corollary 2.8 in [11]). ∎